### v1

# Effect Identification and Unit Categorization in the Multi-Score Regression Discontinuity Design with Application to LED Manufacturing

Abstract
The RDD (regression discontinuity design) is a widely used framework for identification and estimation of causal effects at a cutoff of a single running variable. Practical settings, in particular those encountered in production systems, often involve decision-making defined by multiple thresholds and criteria. Common MRD (multi-score RDD) approaches transform these to a one-dimensional design, to employ identification and estimation results. However, this practice can introduce non-compliant behavior. We develop theoretical tools to identify and reduce some of this “fuzziness” when estimating the cutoff-effect on compliers of sub-rules. We provide a sound definition and categorization of unit behavior types for multi-dimensional cutoff-rules, extending existing categorizations. We identify conditions for the existence and identification of the cutoff-effect on complier in multiple dimensions, and specify when identification remains stable after excluding nevertaker and alwaystaker. Further, we investigate how decomposing cutoff-rules into simpler parts alters the unit behavior. This allows identification and removal of non-compliant units potentially improving estimates. We validate our framework on simulated and real-world data from opto-electronic semiconductor manufacturing. Our empirical results demonstrate the usability for refining production policies. Particularly we show that our approach decreases the estimation variance, highlighting the practical value of the MRD framework in manufacturing.
Keywords: Causal Inference, Machine Learning, Regression Discontinuity, Causal Effect Identification, Multi-Score RDD, Rework, Production Optimization, Phosphor-converted white LEDs
1 Introduction
In Economics and Management – in particular Marketing, Finance, Social sciences, and many other subfields – decision-making often hinges on multiple criteria. Examples include targeted marketing, education policies, political advertising.
The complexity of multiple criteria has led to the development of formal frameworks such as Multi-Criteria Decision Analysis, which structure these trade-offs and provide methods for ranking alternatives. Yet, as decisions become embedded in policy environments, a new challenge emerges: understanding the actual impact of those decisions. Here, the focus shifts from choosing among options to rigorously evaluating outcomes. This is where causal inference enters, and with it, designs like Regression Discontinuity Design (RDD).
RDD has emerged as a powerful quasi-experimental tool for identifying causal effects across diverse fields including economics (Hartmann et al., 2011; Card et al., 2015; Flammer, 2015; Calvo et al., 2019), public policy (Lee, 2008) and the social sciences (Angrist and Lavy, 1999), whenever a policy is in place that leads to a jump in treatment probability at a certain cutoff. Its appeal lies in its ability to deliver credible causal estimates under minimal assumptions. Particularly, RDD can be applied even if typical causal inference assumptions such as unconfoundedness Rubin (1974) or positivity (Austin, 2011) do not hold.
The classic RDD utilizes discontinuities in a single treatment assignment rule that is based on an observed running or score variable. When correctly specified, this setting enables identification of average treatment effects for units near the cutoff. Thus, under mild conditions, it enables to counter-factually reason about the decision policy in place in a small neighborhood around the boundary (Cattaneo et al., 2019; Imbens and Lemieux, 2008; Lee and Lemieux, 2010).
As motivated above, in many real-world contexts – particularly in the scope of industrial, operational, or engineering systems – treatment decisions are typically based on multiple criteria (Sabaei et al., 2015). The logical extension of the classical RDD to multi-score RDD (MRD), in which treatment decisions are taken under a joint logic of multiple score variables satisfying a threshold condition has been proposed by Papay et al. (2011). Influential surveys that propose estimators and empirically study them, include Reardon and Robinson (2012), Wong et al. (2013) and Porter et al. (2017). Particularly in test score-based (An et al., 2024) and geographical (Keele and Titiunik, 2015) applications MRD has been successfully applied in many empirical studies. However, identification strategies of MRD effects for multiple cutoffs in an arbitrarily connected logical rule remains unclear in the literature.
This paper addresses the identification and estimation challenges that arise in the context of multi-score RDD. We develop new identification results tailored for multi-score RDD settings. Specifically, we provide a formal framework that defines and categorizes behavioral types (“complier”, “alwaystaker”, “nevertaker”, “defier”) of units and addresses the inheritance of these types when changing the treatment assignment. We investigate general boolean-type cutoff rules, such as "AND"-type and "OR"-type rules, commonly observed in MRD environments and analyze how their properties influence local identification. Further, we show how the complier effect can be efficiently identified using these categories. We then showcase our results on a dataset of a light-emitting diode (LED) production facility, where each batch undergoes evaluation based on several metrics – such as mean color point measurements or share of units per lot that fulfill a certain quality specification.
Our application of the proposed strategies to data from the LED manufacturing industry demonstrates the utility of our framework in a real-world problem, where treatment assignment is based on multiple quality indicators. We combine our framework with recently proposed estimators that use ML-based adjustments to further reduce variance in estimation (Noack et al., 2024). The estimation of the causal effect of a threshold-based production guides the optimal tuning of the decision threshold to improve manufacturing outcomes. Specifically, managers need to ascertain how marginal adjustments to these thresholds affect outcomes like yield rates and thus the overall production efficiency. We complement our empirical results with a simulated study using a semi-synthetic clone of the production environment, highlighting the value of multi-score RDD for counterfactual analysis and manufacturing policy design.
This study contributes to the operations research literature by bridging causal machine learning advancements with practical decision-making in manufacturing systems. Our theoretical results on effect identification in MRD open new paths for its application in decision making contexts. By accommodating complex, multi-criteria assignment mechanisms, our framework aligns with the realities of operational decision-making, offering tools for more informed and effective policy design.
Our paper is organized as follows. Section 2 reviews related literature on RDD and its extensions to multi-score settings. Section 3 presents our theoretical framework for identification under general Boolean assignment rules. Section 4 details our empirical application in the scope of opto-electronic semiconductor manufacturing. Section 5 provides empirical results of our proposed identification strategies in both synthetic and semi-synthetic environments. Section 6 discusses results from a real data application, highlighting the implications for threshold calibration and production efficiency. Section 7 concludes the paper.
2 Related Literature
The Regression Discontinuity Design dates back to a study on scholarship programs by Thistlethwaite and Campbell (1960). In the following decades, multiple studies employing this quasi-experimental approach gained popularity in the social sciences, such as Angrist and Lavy (1999) and (Black, 1999). Building on this pioneering empirical work, Hahn et al. (2001) first formalized the RD design within the potential-outcomes framework and delivered non-parametric identification and estimation results for both sharp and fuzzy designs, firmly establishing RD as a core tool of modern econometrics.
Impactful surveys including Imbens and Lemieux (2008), Lee and Lemieux (2010) and Cattaneo et al. (2019), who synthesize clear guidance on estimation choices for the local linear regression that serves as estimator, such as bandwidth selection, polynomial order, and diagnostic checks. There were multiple key theoretical contributions recently. Calonico et al. (2014) developed bias-corrected point estimates and robust confidence intervals that improve coverage accuracy and have become the empirical default for RD inference. Calonico et al. (2019b) study the addition of covariates to the local regression in order to reduce the estimator’s variance. Calonico et al. (2019a) study the optimal bandwidth choice for robust RDD estimators. Further extensions that include ML-based covariate adjustments in order to reduce variance in estimation have brought RDD into the toolbox of causal ML researchers (Kreiss and Rothe, 2022; Arai et al., 2025; Noack et al., 2024).
The idea of extending RDD to a multi-score setting dates back to a study from Papay et al. (2011) showing that causal effects can be identified along a two-dimensional cutoff frontier when program eligibility depends jointly on multiple scores. Reardon and Robinson (2012) first systematically collected a set of MRD identification strategies when the running variables jointly cross their cut-offs as a multidimensional “frontier”, translating to an “AND”-connection of rules. They develop both frontier-projection and distance-based estimators, and highlight the steep data-density and bandwidth challenges that arise as the dimensionality of the assignment space grows. Stepping from conceptual proposals to systematic evaluation, Wong et al. (2013) performed extensive simulation studies and demonstrated that estimators which weight observations along the full cutoff frontier achieve lower bias and higher efficiency than naive one-dimension-at-a-time approaches, highlighting the need to exploit the joint assignment rule in practice. In another extensive study, Porter et al. (2017) derive similar findings and give practical advice for MRD settings. An important special case, the geographic RD, tailored to evaluate policies along irregular (state) borders was introduced by Keele and Titiunik (2015). Recently, also more theoretically driven work appeared. Choi and Lee (2018) allow for partial treatment effects when only some running variables cross their cut-offs and introduce thin-plate-spline estimators that flexibly recover these heterogeneous impacts along a high-dimensional frontier. Choi and Lee (2023), building upon earlier work by Imbens and Zajonc (2009), delivered insights on two-dimensional MRD with limited compliance and thus first deriving estimation implications for complier effects in fuzzy MRD.
Recent work has also introduced more advanced MRD estimators based on a minimax approach (Imbens and Wager, 2019), on decision-tree methods (Liu and Qi, 2024), and on fully multivariate local-polynomial estimators (Sawada et al., 2025). Finally, Cattaneo et al. (2025) discuss the statistical properties of important MRD estimators in more detail, giving practical recommendations.
Recent work has proposed the application of the regression discontinuity design in business decision making (Hünermund et al., 2021) or in production operations management (Ho et al., 2017; Mithas et al., 2022). Calvo et al. (2019) use a one-dimensional fuzzy RD design to estimate the effect of procurement officers on budget overruns and delays in public infrastructure projects. Further applications include Leung et al. (2020) who study the effects of the labor unionization of firms on the operating performance of their suppliers. RDD has also been employed to assess the impact of regulation changes on compliance in the car industry (Hu et al., 2021). However, to the best of our knowledge, there are no direct applications of RDD nor MRD to production policies.
This paper contributes to the literature in two main ways: Firstly, we derive theoretical results yielding new results on unit categorization and the identification in MRD that are not limited to, but particularly useful, when dealing with MRD in complex, multi-dimensional decision making. Secondly, we apply our theoretical results in a real-data application, providing a framework that supports the applicability of MRD in operations policy making by deriving implications on a real-world decision policy.
3 General Identification Strategies in Multi-score RDD
In this section, we first introduce the basics of the (Multi-Score) RDD design. Having established this, we give a formal definition of common behavior types of units (e.g., complier, defier) in multi-score, two-stage decision settings that employ cutoff-rules for the initial treatment assignment. We derive an identification result using these unit categories. In particular, we show that under certain assumptions, the identification does not depend on subsets of unit types with constant response. Additionally, we derive rules for the inheritance of unit types as the assignment rules get more complex.
3.1 Regression Discontinuity Design
Recent sources rely on the groundwork of Hahn et al. (2001), which provides a formal justification for effect identification and estimation using RDD. RDD consists of three key ingredients: A score $X$ that defines a certain characteristic of the units, a cutoff $c$ that splits the support of the score into two groups and a treatment $D$, which is assigned to one of the groups based on their score and the cutoff, $D_{i}=\operatorname{\mathbbm{1}}[X_{i}\geq c]$ (Cattaneo et al., 2019).
Typically, a researcher is interested on the average treatment effect at the cutoff $c$, namely $\mathbb{E}[Y(1)-Y(0)\,|\,X=c]$ with $Y(1)$ and $Y(0)$ being the potential outcomes of individuals according to Rubin (2005). The identification of the effect relies on the following key assumption:
Assumption 1 (Continuity).
The conditional mean of the potential outcomes $\mathbb{E}[Y_{i}(d)\mid X_{i}=x]$ for $d\in\{0,1\}$ is continuous at the cutoff level $c$.
This assumption implies that units that receive very similar score on opposite sides of the cutoff are comparable to each other in all relevant aspects (Hahn et al., 2001). To ensure that there is no selection on gains, we have to assure that units around the threshold are not able to perfectly manipulate their score, formally, $\mathbb{E}[Y(1)-Y(0)\mid D,\,X=x]=\mathbb{E}[Y(1)-Y(0)\mid X=x]$ has to hold for $x\in(c-\varepsilon,c+\varepsilon)$ and small $\varepsilon>0$ (Chernozhukov et al., 2024).
Under these relatively mild assumptions, RDD provides inference around the threshold as credible as in a randomized experiment Lee (2008). The average treatment effect at the threshold
$\tau_{0}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid X_{i}=c]$ (1)
can be identified as
$\tau_{0}=\lim_{x\to c^{+}}\mathbb{E}[Y_{i}\mid X_{i}=x]-\lim_{x\to c^{-}}\mathbb{E}[Y_{i}\mid X_{i}=x]$ (2)
Hahn et al. (2001).
3.2 Problem description
While standard RDD is suitable for analyzing effects of a single score variable at a cutoff, real-world treatment decisions often involve multiple score variables. For example, in the case when treatment $T_{i}=1$ is assigned to the unit $i$ only if both criteria $X_{1,i}>c_{1}$ and $X_{2,i}>c_{2}$ are met. Oftentimes, such multi-score cases are reduced to a standard RDD setting by transforming the involved variables into a single score. The identification and estimation of the MRD effect can then be justified by applying theoretical results from the one-dimensional case. One drawback, however, is the reduced interpretability of the transformed score. Thus, the question arises under which precise conditions a treatment effect is existent and identifiable in a multi-score setting. In particular, a well-defined multidimensional cutoff effect requires a certain independence of the directions from which the cutoff is approached.
In applications, one is usually interested in optimizing the treatment assignment $T_{i}$ to improve the outcome $Y_{i}$. Critical systems might only allow for gradual optimization of $T_{i}$. In such cases, a part of the assignment rule $T_{i}$, e.g., $H_{i}\coloneqq\operatorname{\mathbbm{1}}[X_{2,i}>c_{2}]$, is accepted as given, while the remaining part, e.g., $G_{i}\coloneqq\operatorname{\mathbbm{1}}[X_{1,i}>c_{1}]$, is subjected to optimization. Thus, the effect of a subrule $G$ of $T$ at its cutoff can provide insights on how to improve $T$. In setups where the treatment assignment $T$ is seen as a mere recommendation, the actual treatment $D$ is allowed to deviate from it. In RDD, this case is usually referred to as “fuzzy”. All in all, this results in a hierarchy $G\to T\to D$ of decision rules. Each transition potentially obfuscates the assessment of $G$ – in our case, the assessment of $c_{1}$. The reason is clear: There might be units that do not comply with $G$ at each level of the hierarchy. On the other hand, knowledge of the compliance behavior of a unit $i$ with respect to the rules (G,D)(G,\,D) would allow factoring out these obfuscations. This is usually taken care of using a “fuzzy” RDD estimator that adjusts the “sharp” effect estimate by incorporating an estimate of the treatment probability $\Pr(D_{i}=1\,|\,G_{i})$, effectively estimating a complier effect.
We, on the other hand, categorize the unit behavior with respect to the assessed rules (G,D)(G,\,D) into complier, defier, nevertaker, alwaystaker and indecisive items. Using this categorization, we derive assumptions for the identification of a complier effect regarding (G,D)(G,\,D). We investigate conditions under which non-change units (e.g., nevertaker), can be disregarded avoiding some of the obfuscations. In the example, units with $X_{2,i}\leq c_{2}$ always have $T_{i}=0$ regardless of $G_{i}$ and thus should have no influence on the complier effect of $G$. In certain cases, this resulting sub-setting principle can justify the usage of “sharp” RDD estimators, which are typically more stable than their “fuzzy” counterparts. Furthermore, we show how the unit categorization changes when progressing along the hierarchy from simple to complex rules incorporating knowledge about the in between categorizations.
3.3 Setup
We use following notation $\mathbb{R}^{K}_{>0}\coloneqq\left\{x\in\mathbb{R}^{K}\,\middle|\,\forall\,0\leq j\leq K\,:\,x_{j}>0\right\}$ for vectors with positive components and adapt it in the canonical way for the cross product of other ordered sets and relations (e.g. $\mathbb{R}^{K}_{<y}$ or $\mathbb{N}_{\leq k}$). Further, let $e_{1},\ldots,e_{K}\in\mathbb{R}^{K}$ denote the standard basis and $\left\langle\cdot,\,\cdot\right\rangle$ the standard scalar product in $\mathbb{R}^{K}$. For linear spaces $U$ and $V$ over $\mathbb{R}$ with $U\cap V=\{0\}$ we write $U\oplus V$ for their direct sum.
For each individual $i$, let $X_{i}=(X_{1,i},\,\ldots,\,X_{K,i})\in\mathbb{R}^{K}$ denote the score variables. Further, given $c\in\mathbb{R}^{K}$, let $I_{i}(c)\coloneqq(I_{1,i}(c_{1}),\,\ldots,\,I_{K,i}(c_{K}))$ denote the corresponding indicator variables with $I_{k,i}(c_{k})\coloneqq\operatorname{\mathbbm{1}}[X_{k,i}>c_{k}]$ and let the observed outcome for individual $i$ be $Y_{i}$. We regard the entries $I_{i}(c)$ as boolean variables and allow any composition of AND, OR and negation operations ($\land,\,\lor$$\overline{\square}$) over this set of atoms to form general boolean functions $g(I_{i}(c))$.
Definition 1.
A mapping $T:\mathbb{R}^{K}\to\{0,1\}$ is called a decision rule. We say that a decision rule $T$ is a cutoff rule (on $\mathbb{R}^{K}$) if there exists a boolean mapping $g$ such that $T_{i}=T(X_{i})=g(I_{i})$. That is $T$ explains the individual decisions $T_{i}$ using cutoffs according to $X_{i}$.
With slight abuse of notation, we use $T(c)$, $T(X_{i}\,|\,c)$ and $T_{i}(c)$ to indicate the use of a specific cutoff $c\in\mathbb{R}^{K}$. Note that
$T(X+\epsilon\,|\,c)=T(X\,|\,c-\epsilon)\quad\mbox{and}\quad T(\lambda X\,|\,c)=T\left(X\,\middle|\,\frac{c}{\lambda}\right)$ (3)
holds for $\epsilon\in\mathbb{R}^{K},\lambda\in\mathbb{R}_{>0}$. Thus, without loss of generality, we assume that the cutoff of interest is $c=0$. We suppose that $Y_{i}$ depends in the following way on a cutoff rule $T$ and a general decision rule $D$:
$Y_{i}=Y_{i}(T_{i},\,D_{i})$ $=\bigg{(}Y_{i}(0,\,0)(1-D_{i})+Y_{i}(0,\,1)D_{i}\bigg{)}(1-T_{i})$
$+\bigg{(}Y_{i}(1,\,0)(1-D_{i})+Y_{i}(1,\,1)D_{i}\bigg{)}T_{i},$
where $T$ is the treatment assignment and $D$ is the implemented treatment. Unless otherwise stated, we make no further assumptions on $D$ in the proofs, except that it is a decision-rule.
3.4 Unit Categorization
Given this setup, there are certain groups of individuals that are especially interesting, namely the nevertaker, the alwaystaker, the complier and the defier with respect to the pair (T,D)(T,\,D) or (G,D)(G,\,D), where $G$ is a subrule of $T$. We follow the intuition that a meaningful categorization of the behavior of a unit $i$ is only possible by means of counterfactual reasoning. That is, we need to investigate how imagined changes in the cutoff, or equivalently in the observed score values would have changed the responses $T_{i}$ and $D_{i}$.
[…]
Figure 1: Cutoff rules $D=I_{1}\land I_{2}$ and $T=I_{1}$. The decision boundaries coincide with the coordinate axes. $T$ complies with $D$ given $X_{2}>0$. Otherwise, $D$ is zero independent of all variables that control the behavior of $T$, namely $X_{1}$.
Figure 1 illustrates this idea for $D\coloneqq I_{1}\land I_{2}$ and $T\coloneqq I_{1}$, and shows that only changes $c\in\mathbb{R}^{2}$ that impact $T$ are relevant to categorize the behavior of a unit $i$ regarding (T,D)(T,\,D). Additional variables that only affect $D$ might not be controllable or even known. The following definition formalizes the intuition behind relevant change directions of a cutoff rule.
Definition 2.
Let $T$ be a cutoff rule on $\mathbb{R}^{K}$. Denote with
$S(T)\coloneqq\left\{k\in\mathbb{N}_{\leq K}\,\middle|\,\exists c\in\mathbb{R}^{K},\lambda\in\mathbb{R}:\,T(0\,|\,c+\lambda e_{k})\neq T(0\,|\,c)\right\}$
the support directions of $T$. The support of $T$ is defined as the linear hull over the support directions:
$\operatorname{supp}(T)\coloneqq\left\{\sum_{k\in S(T)}\lambda_{k}e_{k}\,\middle|\,\lambda_{k}\in\mathbb{R},\,k\in S(T)\right\}$
This rather technical definition greatly simplifies the treatment of equivalent cutoff-rules e.g. $T\coloneqq I_{1}\land(I_{2}\lor\overline{I_{2}})$ and $G\coloneqq I_{1}$. Additionally, it allows comparing $T$ and $D$ on their common domain. While above notion of the support of a cutoff rule is motivated by the existence of a change along a single coordinate direction, it is almost trivial to see that an empty support also excludes changes along multiple directions. This justifies the definition.
Lemma 1.
$T$ is constant if and only if $S(T)=\emptyset$, or equivalently, if and only if $\operatorname{supp}(T)=\{0\}$.
From now on, we require that $T$ is not degenerate in the sense that $\operatorname{supp}(T)\neq\{0\}$. For the introductory AND-rule example, $D\coloneqq I_{1}(c)\land I_{2}(c)$ and $T\coloneqq I_{1}(c)$ we have $\operatorname{supp}(T)=\mathbb{R}\times\{0\}$ and $\operatorname{supp}(D)=\mathbb{R}^{2}$. If a unit $i$ complies in the assigned treatment $T$ with the actual $D$, one would require that $T_{i}$ produces the same output as $D_{i}$, even under any hypothetical change of the cutoff $c_{1}$ as indicated in Figure 1. A more refined, but local definition could relax this requirement to “any reasonable changes”. For ease of presentation we stick to the global one. In other words, both rules should be synchronous on the support of $T$, which is the case if $X_{2,i}>c_{2}$. Thus, it is $X_{2,i}$ (or equivalently $c_{2}$) that controls the behavior of $i$. To make this distinction more apparent, it is useful to have a notation for entries $X\in\mathbb{R}^{K}$ that do not affect $T$:
$N^{T}\coloneqq\left\{X\in\mathbb{R}^{K}\,\middle|\,T(X\,|\,c)=T(0\,|\,c)\,\mbox{for all}\,c\in\mathbb{R}^{K}\right\}$
In particular, one can show the following.
Proposition 1.
For each $X\in\mathbb{R}^{K}$, there exists a unique decomposition $X=X^{T}+X^{\perp T}$ with $X^{T}\in\operatorname{supp}(T)$ and $X^{\perp T}\in N^{T}$. The decomposition according to the orthogonal projection $P_{T}(X)\coloneqq\sum_{k\in S(T)}\left\langle X,\,e_{k}\right\rangle e_{k}$ onto $\operatorname{supp}(T)$ satisfies the above properties.
That is, according to Proposition 1 one has the following decomposition
$\mathcal{S}\coloneqq\mathbb{R}^{|S(T)|}\times\mathbb{R}^{K-|S(T)|}\simeq\operatorname{supp}(T)\,\oplus N^{T}=\mathbb{R}^{K}$
of the score space into two distinct parts. The former part captures the decision changes of $T$ and the latter consists of free variables not affecting $T$, but potentially controlling the unit category. This observation allows for a general definition of unit categories.
Definition 3.
Let $i$ be a unit and $X_{i}=X_{i}^{T}+X_{i}^{\perp T}$. Then $i$ is said to be
  1. 1.
    a nevertaker (of $T$ with respect to $D$) iff $D(X_{i}^{\perp T}-c)=0$
  2. 2.
    an alwaystaker (of $T$ with respect to $D$) iff $D(X_{i}^{\perp T}-c)=1$
  3. 3.
    a complier (of $T$ with respect to $D$) iff $T(0\,|\,c)=D(X_{i}^{\perp T}-c)$
  4. 4.
    a defier (of $T$ with respect to $D$) iff $T(0\,|\,c)\neq D(X_{i}^{\perp T}-c)$
for all $c\in\operatorname{supp}(T)$. Let the set of complier, nevertaker, alwaystaker and defier be denoted by $\operatorname{ComP}(T,D)$, $\operatorname{At}(T,D)$, $\operatorname{Nt}(T,D)$ and $\operatorname{DeF}(T,D)$.
Particularly, if $D$ is a cutoff-rule one obtains the following more intuitive equivalencies, which capture the notion of a simultaneous cutoff changes along relevant directions.
Proposition 2.
Let $D$ be a cutoff rule over $\mathbb{R}^{K}$. A unit $i$ is
  1. 1.
    a nevertaker (of $T$ with respect to $D$) iff $D_{i}(c)=0$
  2. 2.
    an alwaystaker (of $T$ with respect to $D$) iff $D_{i}(c)=1$
  3. 3.
    a complier (of $T$ with respect to $D$) iff $T_{i}(c)=D_{i}(c)$
  4. 4.
    a defier (of $T$ with respect to $D$) iff $D_{i}(c)\neq T_{i}(c)$
for all $c\in\operatorname{supp}(T)$.
The terminology established in Definition 3 indeed introduces well-defined categories:
Proposition 3.
The sets $\operatorname{At}(T,\,D),\,\operatorname{Nt}(T,\,D),\,\operatorname{ComP}(T,\,D)$ and $\operatorname{DeF}(T,\,D)$ are pairwise disjoint.
[…]
(a) No matter which change $c$ of $X_{i}$ in the $\operatorname{supp}(T)$ plane we imagine, the responses of $D$ and $T$ are equal. A shift of score $X_{i}$ by $c$ is equivalent to a simultaneous shift of the decision boundaries by $c$, which amounts to a coordinate shift for $D$ and a cutoff shift for $T$ motivating Definition 3.
[…]
(b) Unit categorizations regarding (T,D)(T,\,D). While $\operatorname{supp}(T)$ captures change directions of $T$, the unit behavior is determined by $N^{T}$. The space is divided into (horizontally depicted) areas of alwaystaker, indecisives, complier, indecisives and nevertakers of $T$ with respect to $D$.
Figure 2: General case with $D$ not being a cutoff-rule. The decision boundaries of $T$ and $D$ are indicated with orange and blue dashed lines.
From now on, whenever we assume that $D$ is a cutoff rule, we restrict ourselves to the case in which $D$ has at least as much information for decision-making as $T$. This means that $D$ might depend on $I_{k,i}(0)$ for $k\in S(T)$. In addition, we suppose that $D$ does not depend on $I_{k,i}(c)$ with $c\neq 0$ for $k\in\operatorname{supp}(T)$, effectively restricting to the case where both decision rules have a zero cutoff. This restricts the general case. Suppose that the decision maker $D$ is an opportunist and ignores all the other scores as long as a certain incentive $X_{k}$ exceeds an even higher cutoff $0<c_{k}<X_{k}$. Only then $D$ would comply with $T$.
Even for cutoff rules $D$ with these restrictions, indecisive units are possible. For example let $D\coloneqq(\overline{I}_{1}\land I_{2})$ and $T\coloneqq I_{1}\land I_{2}$. Then $\operatorname{supp}(T)=\mathbb{R}^{2}=\operatorname{supp}(D)$ and for T(0| 0)=D(0| 0)T(0\,|\,0)=D(0\,|\,0) but $0=T(0\,|(-1,1))\neq D(0\,|\,(-1,1))=1$. Thus, $D$ is not constant nor equal to $T$ or $\overline{T}$ on the support of $T$. Given $T$ with $\dim(\operatorname{supp}(T))>1$, one can always construct a cutoff rule $D$ that exhibits indecisive items. At least one has the following:
Proposition 4.
Let $i$ denote an individual. Then $i\in\operatorname{At}(T,D)\cup\operatorname{Nt}(T,D)\cup\operatorname{DeF}(T,D)\cup\operatorname{ComP}(T,D)$ for all cutoff rules $D$ if and only if $\dim(\operatorname{supp}(T))=1$.
Moreover, our definitions of complier, nevertaker and alwaystaker imply the corresponding definitions in Imbens and Lemieux (2008). For this, let $c^{+},c^{-}\in\operatorname{supp}(T)$ be two directions that induce change in $T$, that is:
$\lim_{\lambda\to 0}T(X_{i}\,|\,\lambda c^{+})=1\,\mbox{ and }\,\lim_{\lambda\to 0}T(X_{i}\,|\,\lambda c^{-})=0$
Using the above complier definition:
$T(X_{i}\,|\,c)=T(X_{i}^{T}\,|\,c)=T(0\,|\,c-X_{i}^{T})=D(X_{i}^{\perp T}+X_{i}^{T}-c)=D(X_{i}-c)$
for $c\in\operatorname{supp}(T)$. In particular, $T$ and $D$ coincide in a neighborhood of zero, and thus
$\lim_{\lambda\to 0}D_{i}(\lambda c^{+})=\lim_{\lambda\to 0}D(X_{i}-\lambda c^{+})=1\,\mbox{ and }\,\lim_{\lambda\to 0}D_{i}(\lambda c^{-})=\lim_{\lambda\to 0}D(X_{i}-\lambda c^{-})=0$
where $D_{i}(c)\coloneqq D(X_{i}-c)$ for $c\in\operatorname{supp}(T)$. The corresponding statements for nevertaker and alwaystaker follow analogously. Indeed, by requiring consistency in the limit for any direction, one would arrive at a local definition of the unit categories that is sufficient for the multi-score RDD setting. Before continuing with the identification part, we are going to expand on the AND-rule example.
Example (AND-Rules)
Let $D\coloneqq\bigwedge_{j=1}^{K}I_{j}$ and $T\coloneqq\bigwedge_{j=1}^{k}I_{j}$ for $k\in\mathbb{N}_{<K}$. Then $\operatorname{supp}(T)=\mathbb{R}^{k}\times\{0\}^{K-k}$, and the potential outcomes reduce to
$Y_{i}=Y_{i}(0,0)(1-T_{i})+Y_{i}(1,0)(1-D_{i})T_{i}+Y(1,1)D_{i}$
, dropping one of the mixed terms. Additionally, we have the following unit categorizations:
  1. 1.
    $\operatorname{ComP}(T,\,D)=\{i\,|\,X_{i}\in\mathbb{R}^{k}\times\mathbb{R}^{K-k}_{>0}\}$
    Note that $X_{i}\in\mathbb{R}^{k}\times\mathbb{R}^{K-k}_{>0}$ is equivalent to $\land_{j=k+1}^{K}I_{j,i}(c)$ being one for $c\in\operatorname{supp}(T)$. Thus, the rule $D_{i}$ effectively reduces to $T_{i}$.
  2. 2.
    $\operatorname{At}(T,\,D)=\emptyset$
    Since for each $X_{i}\in\mathbb{R}^{K}$ there exists a $c\in\operatorname{supp}(T)$ such that $D_{i}(c)=0$. For example choose any $c$ with $X_{1,i}\leq c_{1}$ and $c_{j}=0$ for $j\neq 1$.
  3. 3.
    $\operatorname{Nt}(T,\,D)=\{i\,|\,\exists k<j\leq K:\,X_{j,i}\leq 0\}=\{i\,|\,X_{i}\in\mathbb{R}^{k}\times(\mathbb{R}^{K-k}\setminus\mathbb{R}^{K-k}_{>0})\}$
    Note that $i$ being in the set on the right is equivalent to $\land_{j=k+1}^{K}I_{j,i}(c)$ being zero for all $c\in\operatorname{supp}(T)$. This is equivalent to $D_{i}(c)=0$ for all $c\in\operatorname{supp}(T)$, since given an $i$ one can always find a cutoff $c\in\operatorname{supp}(T)$ such that $T_{i}(c)=1$, which implies that one of the indicators $I_{j,i}$ for $j>k$ has to be zero.
  4. 4.
    $\operatorname{DeF}(T,\,D)=\emptyset$
    Note that $T_{i}(c)=0$ implies $D_{i}(c)=0$, and that $c\in\operatorname{supp}(T)$ can always be chosen such that the former is satisfied.
Further examples can be found in the Appendix A.2 showing instances for the remaining unit categories (excluding the indecisive case). We conclude that the introduced definitions are indeed reasonable.
3.5 Effect Identification
Inspired by the work of Hahn et al. (2001); Imbens and Angrist (1994) we are going to use the introduced unit categories to prove an identification theorem for the effect on complier at the cutoff. For this section, we require that the outcome $Y$ does not directly depend on the treatment assignment $T$. Denote the set of all unit categories by
$\mathcal{C}\coloneqq\{\operatorname{ComP}(T,\,D),\,\operatorname{Nt}(T,\,D),\,\operatorname{At}(T,\,D),\,\operatorname{DeF}(T,\,D),\,\operatorname{Ind}(T,\,D)\}$
and the set of non-change categories by
$\mathcal{C}^{0}\coloneqq\{\operatorname{Nt}(T,\,D),\,\operatorname{At}(T,\,D)\}.$
We assume that the categorization of a unit is independent of the support part of $T$ in a neighborhood of the cutoff, that is:
Assumption 2.
There exists an $\epsilon>0$ such that $\Pr\left(i\in\operatorname{Cat}\,\middle|\,X_{i}^{T}=x\right)=\Pr\left(i\in\operatorname{Cat}\,\middle|\,X_{i}^{T}=0\right)$ for $\|x\|\leq\epsilon$ and $\operatorname{Cat}\in\mathcal{C}$.
This requirement relates to the independence assumptions used in Imbens and Angrist (1994). Using Assumption 2 and further assuming $\Pr\left(i\in\operatorname{Cat}\,\middle|\,X^{T}_{i}=0\right)>0$ one has
$\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=x\right)=\sum_{\operatorname{Cat}\in\mathcal{C}}\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=x,\,i\in\operatorname{Cat}\right)\Pr\left(i\in\operatorname{Cat}\,\middle|\,X_{i}^{T}=0\right)$
and thus
$\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=x^{+}\right)-\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=x^{-}\right)$ $=$
$\sum_{\operatorname{Cat}\in\mathcal{C}}\bigg{(}\mathbb{E}\left(Y_{i}\,\middle|X_{i}^{T}=x^{+},\,i\in\operatorname{Cat}\right)$ $-\mathbb{E}\left(Y_{i}\,\middle|X_{i}^{T}=x^{-},\,i\in\operatorname{Cat}\right)\bigg{)}\Pr\left(i\in\operatorname{Cat}\,\middle|\,X_{i}^{T}=0\right)$
We further rely on the following local continuity assumption, which is a variation of the standard continuity assumption (Assumption 1) adapted to the unit categories:
Assumption 3.
There exists an $\epsilon>0$ such that $x\mapsto\mathbb{E}(Y_{i}(d)\,|\,X_{i}^{T}=x,\,i\in\operatorname{Cat})$ is continuous for $\|x\|\leq\epsilon$, $d\in\{0,1\}$ and $\operatorname{Cat}\in\mathcal{C}$.
This assumption can be weakened by assuming only directional continuity, which would render the effect dependent on the chosen directions. Note that
$E\left(Y_{i}\,\middle|\,X_{i}^{T}=x^{\pm},\,i\in\operatorname{At}\right)=E\left(Y_{i}(1)\,\middle|\,X_{i}^{T}=x^{\pm},\,i\in\operatorname{At}\right)$
and
$E\left(Y_{i}\,\middle|\,X_{i}^{T}=x^{\pm},\,i\in\operatorname{Nt}\right)=E\left(Y_{i}(0)\,\middle|\,X_{i}^{T}=x^{\pm},\,i\in\operatorname{Nt}\right)$
holds for the non-change unit categories. Together with Assumption 3 one derives:
$\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{+}\right)$ $-\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{-}\right)=$ (4)
$\sum_{\operatorname{Cat}\in\mathcal{C}\setminus\mathcal{C}^{0}}\bigg{(}\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|X_{i}^{T}=\lambda x^{+},\,i\in\operatorname{Cat}\right)$ $-\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|X_{i}^{T}=\lambda x^{-},\,i\in\operatorname{Cat}\right)\bigg{)}\cdot$
$\cdot\Pr\left(i\in\operatorname{Cat}\,\middle|\,X_{i}^{T}=0\right)$
Two more assumptions are required to make further use of the continuity. First, we deny the existence of indecisive items, since this category does not allow structured conclusions about $D$ based on knowledge of $T$. That is, this category does not separate the potential outcomes $Y_{i}(0)$ and $Y_{i}(1)$.
Assumption 4.
$\operatorname{Ind}(T,\,D)=\emptyset$
Second, we assume that the directions $x^{+},\,x^{-}\in\operatorname{supp}(T)$ along which we aim to estimate the complier effect induce a change in $T$.
Assumption 5.
$1=\lim_{\lambda\to 0}T\left(\lambda x^{+}\,\middle|\,0\right)\neq\lim_{\lambda\to 0}T\left(\lambda x^{-}\,\middle|\,0\right)=0$
This assumption is implicit in one-dimensional RDD designs and imposes no real restriction in practice, as $T$ is generally assumed to be known. With this, we know how $D_{i}$ behaves for complier and defier when approaching from $x^{+}$ and $x^{-}$ direction. That is:
$\lim_{\lambda\to 0}\Pr\left(D_{i}=1\,|\,X_{i}^{T}=\lambda x^{+},\,i\in\operatorname{ComP}\right)=1\,\mbox{ and }\,\lim_{\lambda\to 0}\Pr\left(D_{i}=1\,|\,X_{i}^{T}=\lambda x^{-},\,i\in\operatorname{ComP}\right)=0,$
as well as
$\lim_{\lambda\to 0}\Pr\left(D_{i}=1\,|\,X_{i}^{T}=\lambda x^{+},\,i\in\operatorname{DeF}\right)=0\,\mbox{ and }\,\lim_{\lambda\to 0}\Pr\left(D_{i}=1\,|\,X_{i}^{T}=\lambda x^{-},\,i\in\operatorname{DeF}\right)=1.$
Thus, we can apply Assumption 3 to these two remaining categories on the right side of Equation 4 as well:
Theorem 1.
Let Assumptions 2, 3, 4 and 5 hold. Then the complier effect at the cutoff is identified as
$\mathbb{E}\left(Y_{i}(1)\,\middle|\,X_{i}^{T}=0,\,i\in\operatorname{ComP}\right)$ $-\mathbb{E}\left(Y_{i}(0)\,\middle|\,X_{i}^{T}=0,\,i\in\operatorname{ComP}\right)=$ (5)
$\frac{1}{\Pr\left(i\in\operatorname{ComP}\,\middle|\,X_{i}^{T}=0\right)}$ $\left(\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{+}\right)-\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{-}\right)\right)-C$
with $C$ being the correction term for defier:
$C\coloneqq\frac{\Pr\left(i\in\operatorname{DeF}\,\middle|\,X_{i}^{T}=0\right)}{\Pr\left(i\in\operatorname{ComP}\,\middle|\,X_{i}^{T}=0\right)}\bigg{(}\mathbb{E}\left(Y_{i}(0)\,\middle|\,X_{i}^{T}=0,\,i\in\operatorname{DeF}\right)-\mathbb{E}\left(Y_{i}(1)\,\middle|\,X_{i}^{T}=0,\,i\in\operatorname{DeF}\right)\bigg{)}.$
This identification result allows for two immediate conclusions. First, one is free to choose among the directions $x^{+}$ and $x^{-}$ satisfying Assumption 5. Second, the proof suggests that dropping subsets $\Omega\subset\operatorname{Nt}\cup\operatorname{At}$ does not affect identification, as long as doing so does not violate Assumptions 3 and 2.
Following this thought we now investigate how dropping units in $\operatorname{At}\cup\operatorname{Nt}$ affects the above identification result. For ease of presentation, we assume that there do not exist any defier at the cutoff. As a consequence, the term $C$ in Equation 5 is equal to zero. Now let $\Omega\subset\operatorname{At}\cup\operatorname{Nt}$. Then $\Pr\left(i\in\operatorname{ComP},\,i\in\Omega\,\middle|\,X_{i}^{T}=0\right)=0$ and thus one has
$\Pr\left(i\in\operatorname{ComP}\,\middle|\,X_{i}^{T}=0\right)=\Pr\left(i\in\operatorname{ComP}\,\middle|\,X_{i}^{T}=0,\,i\notin\Omega\right)\Pr\left(i\notin\Omega\,\middle|\,X_{i}^{T}=0\right)$ (6)
for the denominator in Equation 5. For the nominator,
$\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{\pm},\,i\in\Omega\right)$ $=\mathbb{E}\left(Y_{i}(0)\,\middle|\,X_{i}^{T}=\lambda x^{\pm},\,i\in\Omega\cap\operatorname{Nt}\right)\Pr\left(i\in\operatorname{Nt}\,\middle|\,X_{i}^{T}=\lambda x^{\pm},\,i\in\Omega\right)$
$+\mathbb{E}\left(Y_{i}(1)\,\middle|\,X_{i}^{T}=\lambda x^{\pm},\,i\in\Omega\cap\operatorname{At}\right)\Pr\left(i\in\operatorname{At}\,\middle|\,X_{i}^{T}=\lambda x^{\pm},\,i\in\Omega\right)$
holds. Requiring Assumption 3 and Assumption 2 to hold when conditioning on $\Omega\cap\operatorname{Nt}$ and $\Omega\cap\operatorname{At}$ (instead of $\operatorname{Nt}$ and $\operatorname{At}$), we get:
$\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{+},\,i\in\Omega\right)$ $-\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{-},i\in\Omega\right)=0.$ (7)
Since
$\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{\pm}\right)$ $=\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{\pm},\,i\in\Omega\right)\Pr\left(i\in\Omega\,\middle|\,X_{i}^{T}=\lambda x^{\pm}\right)$
$+\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{\pm},\,i\notin\Omega\right)\Pr\left(i\notin\Omega\,\middle|\,X_{i}^{T}=\lambda x^{\pm}\right)$
holds, we require an assumption similar to Assumption 2 for $\Pr\left(i\in\Omega\,\middle|\,X_{i}^{T}=\lambda x^{\pm}\right)$, in order to get
$\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{+}\right)$ $-\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{-}\right)$ (8)
$=\left(\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{+},\,i\notin\Omega\right)-\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{-},\,i\notin\Omega\right)\right)\cdot$
$\cdot\Pr\left(i\notin\Omega\,\middle|\,X_{i}^{T}=0\right)$
using Equation 7. Combining Equations 8 and 6, we obtain the following result:
Theorem 2.
Let Assumptions 2, 3, 4 and 5 hold, and $\Pr(i\in\operatorname{DeF}\,|\,X_{i}^{T}=0)=0$. Further, let $\Omega\subset\operatorname{Nt}\cup\operatorname{At}$ such that
  1. 1.
    there exists an $\epsilon>0$ such that the functions
    $x\mapsto\mathbb{E}\left(Y_{i}(0)\,\middle|\,X_{i}^{T}=x,\,i\in\Omega\cap\operatorname{Nt}\right)$
    and
    $x\mapsto\mathbb{E}\left(Y_{i}(1)\,\middle|\,X_{i}^{T}=x,\,i\in\Omega\cap\operatorname{At}\right)$
    are continuous for $\|x\|<\epsilon$.
  2. 2.
    there exists an $\epsilon>0$ such that
    $\Pr\left(i\in\operatorname{Nt}\,\middle|\,X_{i}^{T}=x,\,i\in\Omega\right)=\Pr\left(i\in\operatorname{Nt}\,\middle|\,X_{i}^{T}=0,\,i\in\Omega\right)$
    and
    $\Pr\left(i\in\operatorname{At}\,\middle|\,X_{i}^{T}=x,\,i\in\Omega\right)=\Pr\left(i\in\operatorname{At}\,\middle|\,X_{i}^{T}=0,\,i\in\Omega\right)$
    as well as
    $\Pr\left(i\in\Omega\,\middle|\,X_{i}^{T}=x\right)=\Pr\left(i\in\Omega\,\middle|\,X_{i}^{T}=0\right)$
    for $\|x\|<\epsilon$.
Then
$\mathbb{E}\left(Y_{i}(1)\,\middle|\,X_{i}^{T}=0,\,i\in\operatorname{ComP}\right)$ $-\mathbb{E}\left(Y_{i}(0)\,\middle|\,X_{i}^{T}=0,\,i\in\operatorname{ComP}\right)=$
$\frac{1}{\Pr\left(i\in\operatorname{ComP}\,\middle|\,X_{i}^{T}=0,\,i\notin\Omega\right)}$ $\left(\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{+},\,i\notin\Omega\right)-\lim_{\lambda\to 0}\mathbb{E}\left(Y_{i}\,\middle|\,X_{i}^{T}=\lambda x^{-},i\notin\Omega\right)\right)$
holds.
We call estimates of the complier effect of $T$ when removing $\Omega$ the sub-set complier effect of $T$ excluding $\Omega$. Theorem 2 provides sufficient conditions when both effects (the complier effect as identified in Theorem 1 and the sub-set complier effect) are equal.
3.6 Inheritance
In this section, we investigate how the unit categorization changes when altering the treatment assignment $T$, e.g. replacing it with a subrule $G$ of $T$. The goal is to avoid using knowledge of potential unobserved parts of the involved decision rules when identifying the unit behavior. Particularly, we are going to relate the unit behavior of complex rules to simpler subrules. First we investigate how the categories behave under the negation operation:
Proposition 5.
The following duality statements hold:
  1. 1.
    $\operatorname{At}(T,\,D)=\operatorname{Nt}(T,\,\overline{D})$
  2. 2.
    $\operatorname{At}(T,\,D)=\operatorname{At}(\overline{T},\,D)$ and $\operatorname{Nt}(T,\,D)=\operatorname{Nt}(\overline{T},\,D)$
  3. 3.
    $\operatorname{Ind}(T,\,D)=\operatorname{Ind}(\overline{T},\,D)=\operatorname{Ind}(T,\,\overline{D})$
  4. 4.
    $\operatorname{ComP}(T,\,D)=\operatorname{DeF}(\overline{T},\,D)=\operatorname{DeF}(T,\,\overline{D})$
Indecisive items remain stable under negation of both decision rules whereas the non-change units are stable under negation of $T$ but flip when negating $D$. This is in contrast to complier/defier units that always flip. Next, we investigate the stability of the non-change categories under sub-rules. In general one has the following bounds:
Proposition 6.
Let $G$ be a cutoff rule on $\mathbb{R}^{K}$ with $\operatorname{supp}(G)\subset\operatorname{supp}(T)$.
  1. 1.
    $\operatorname{At}(T,\,D)\subset\operatorname{At}(G,\,D)$
  2. 2.
    $\operatorname{Nt}(T,\,D)\subset\operatorname{Nt}(G,\,D)$
To obtain similar results for the other categories, assumptions about the relation between $G$ and $T$ are required. The next statement generalizes the AND-rule and OR-rule examples (Example 3.4 and Example A.2).
Proposition 7.
Let G,HG,\,H be cutoff rules on $\mathbb{R}^{K}$ with $\operatorname{supp}(G)\neq\{0\}$. Further, let $\operatorname{supp}(T)=\operatorname{supp}(G)\oplus\operatorname{supp}(H)$.
  1. 1.
    If $T=G\land H$ then:
    1. (a)
      $\operatorname{ComP}(G,\,T)=\{i\,|\,H(X_{i}^{H}\,|\,0)=1\}$
    2. (b)
      $\operatorname{Nt}(G,\,T)=\{i\,|\,H(X_{i}^{H}\,|\,0)=0\}$
    3. (c)
      Every unit $i$ is either a complier or a nevertaker of $G$ with respect to $T$.
  2. 2.
    If $T=G\lor H$ then:
    1. (a)
      $\operatorname{ComP}(G,\,T)=\{i\,|\,H(X_{i}^{H}\,|\,0)=0\}$
    2. (b)
      $\operatorname{At}(G,\,T)=\{i\,|\,H(X_{i}^{H}\,|\,0)=1\}$
    3. (c)
      Every unit $i$ is either a complier or a alwaystaker of $G$ with respect to $T$.
In the above setting, if the sub-rule $H$ is known, we are able to identify the unit categories regarding (G,T)(G,\,T), even if $G$ is not known. Moreover, using Proposition 7 we conclude that the introduction of new score variables in cutoff-rules does not introduce any defier or indecisives. This is not too surprising since, by definition, negation is required to obtain defier (see also Proposition 5). The previous statement can be extended to the case of a general decision rule $D$ setting as follows:
Proposition 8.
Let G,HG,\,H be cutoff rules on $\mathbb{R}^{K}$ with $\operatorname{supp}(T)=\operatorname{supp}(G)\oplus\operatorname{supp}(H)$ and $\operatorname{supp}(G)\neq\{0\}\neq\operatorname{supp}(H)$.
  1. 1.
    If $T=G\land H$ then:
    1. (a)
      $\operatorname{ComP}(G,\,T)\cap\operatorname{ComP}(T,\,D)\subset\operatorname{ComP}(G,\,D)$
    2. (b)
      $\operatorname{Nt}(G,\,T)\cap\operatorname{ComP}(T,\,D)\subset\operatorname{Nt}(G,\,D)$
    3. (c)
      $\operatorname{ComP}(T,\,D)\subset\operatorname{ComP}(G,\,D)\cup\operatorname{Nt}(G,\,D)$
  2. 2.
    If $T=G\lor H$ then:
    1. (a)
      $\operatorname{ComP}(G,\,T)\cap\operatorname{ComP}(T,\,D)\subset\operatorname{ComP}(G,\,D)$
    2. (b)
      $\operatorname{At}(G,\,T)\cap\operatorname{ComP}(T,\,D)\subset\operatorname{At}(G,\,D)$
    3. (c)
      $\operatorname{ComP}(T,\,D)\subset\operatorname{ComP}(G,\,D)\cup\operatorname{At}(G,\,D)$
Thus, complier of $T\coloneqq G\,\square\,H$ with respect to $D$ are bounded by complier and non-change units of the simpler rule $G$ by Proposition 8 2c and 8 1c. This equality does not hold in general. For this set, let $D\coloneqq(I_{1}\land I_{2})\lor I_{3}$, $T\coloneqq I_{1}\land I_{3}$ and $G\coloneqq I_{1}$. Then
$\emptyset=\operatorname{ComP}(T,\,D)\neq\operatorname{ComP}(G,\,D)=\{i\,|\,X_{3,i}\leq 0,\,X_{2,i}>0\}$
holds. Further, Proposition 8 2a, 8 1a and 8 2b, 8 1b resemble a factorization rule. Loosely speaking, the in between part $T$ of the rule hierarchy (G,T)(G,\,T) and (T,D)(T,\,D) factors out, becoming (G,D)(G,\,D). Using Proposition 8 and 5, we can immediately draw conclusions about the defier:
Proposition 9.
Let G,HG,\,H be cutoff rules on $\mathbb{R}^{K}$ with $\operatorname{supp}(T)=\operatorname{supp}(G)\oplus\operatorname{supp}(H)$ and $\operatorname{supp}(G)\neq\{0\}\neq\operatorname{supp}(H)$.
  1. 1.
    If $T=G\land H$ then:
    1. (a)
      $\operatorname{ComP}(G,\,T)\cap\operatorname{DeF}(T,\,D)\subset\operatorname{DeF}(G,\,D)$
    2. (b)
      $\operatorname{Nt}(G,\,T)\cap\operatorname{DeF}(T,\,D)\subset\operatorname{At}(G,\,D)$
    3. (c)
      $\operatorname{DeF}(T,\,D)\subset\operatorname{DeF}(G,\,D)\cup\operatorname{At}(G,\,D)$
  2. 2.
    If $T=G\lor H$ then:
    1. (a)
      $\operatorname{ComP}(G,\,T)\cap\operatorname{DeF}(T,\,D)\subset\operatorname{DeF}(G,\,D)$
    2. (b)
      $\operatorname{At}(G,\,T)\cap\operatorname{DeF}(T,\,D)\subset\operatorname{Nt}(G,\,D)$
    3. (c)
      $\operatorname{DeF}(T,\,D)\subset\operatorname{DeF}(G,\,D)\cup\operatorname{Nt}(G,\,D)$
Combining both statements, we note that if a unit complies regarding (T,D)(T,\,D) the categorization regarding (G,T)(G,\,T) persists for (G,D)(G,\,D). Figure 3 summarizes the above findings for the AND-case. Further, according to Proposition 7, in the absence of indecisive items these transitions are exhaustive, since any unit is either a complier or nevertaker of $G$ with respect to $T$.
[…]
Figure 3: All possible category changes when transitioning from the complex rule $T=G\land H$ to the simpler rule $G$ given that $\operatorname{supp}(G)\oplus\operatorname{supp}(H)=\operatorname{supp}(T)$ and that there are no indecisives. The image summarizes Propositions 6, 7, 8 and 9. The OR-case can be summarized accordingly.
4 Empirical Application
In this section, we apply our theoretical results to a real-world decision-making scenario in opto-electronic semiconductor manufacturing. In particular, we investigate the effect of rework decisions during the color-conversion process in white light emitting diode (wLED) production on the product yield. After a brief description of the application, we derive some modeling implications in accordance to the observed data. This leads to an algorithmic description of the process, which is used as a data-generating process (DGP) complementing the empirical study with simulated runs based on artificial generated data that mimics the real-world case.
4.1 Rework
The phosphor conversion is a crucial step during production of wLEDs. In order to obtain a white color, several layers of phosphor are applied to a blue light emitting ground substrate, shifting the perceived color along a conversion curve from the blue region to the white region. During this conversion step, a production lot, consisting of $784$ individual wLEDs, is processed according to the product recipe. Each wLED in the lot undergoes the same production procedure, resulting in an identical stack of phosphor layers. The overall goal is to reach the specified target color after production is finished. To achieve this, intermediate color measurements are taken after the conversion step assessing proximity to the conversion target, for selected wLEDs in the lot. The processing steps after color conversion are adjusted so that a maximum of wLEDs reach the target color, as long as they are already close to it. In order to ensure a high yield, that is a high share of wLEDs in the lot that reach the color target, a rework decision is made using these intermediate measurements. That is, if the conversion target is not reached, another correction layer of phosphor is applied to the lot. For details on the conversion process, we refer to Cho et al. (2017) and Schwarz et al. (2024).
The specific decision is made based on two scores: The distance score, $X_{D}$, measures the distance of the mean color point $C_{1}=(C_{x},\,C_{y})$ measured after the regular conversion to the conversion target $C_{T}:=(C_{T,x},\,C_{T,y})$. The yield improvement score, $X_{Y}$, is a relative measure that evaluates the quality distribution of individual chips in a lot by calculating a hypothetical scenario in which the target is ideally met by the mean of the lot. If there is high variability in within a lot, rework treatment might have an overall negative effect on the yield even if the distance score suggests a treatment. Figure 4 depicts the decision criteria in more detail.
Figure 4: Left: $X_{D}$ is defined as the distance between the current mean color point $C_{1}$ and the target point $C_{P}$, which is the closest the lot can technically get to the initial target $C_{T}$. The slope of the dashed curve is given by a physical process and thus it is not possible to reach $C_{T}$ once the curve deviates. Right: $X_{Y}$ evaluates the expected improvement by calculating the share of in-specification chips in the lot. This is done by moving the current distribution of color points to the target.
The treatment is assigned according to the cutoff rule $T=I_{D}\land I_{Y}$ in order to maximize the outcome $Y$, which is the percentage of chips found that have reached the desired target after the production has finished.
However, an examination of the empirical data reveals that the operators responsible for performing the actual treatment $D$ do not always comply with $T$ (see Figure 5). We assume that this discrepancy is due to an informational advantage regarding the $X_{Y}$ score that the final decision makers (i.e., human operators) $D$ might have and which they use to override $T$, in order to avoid possible degradation due to the additional rework layering.
We formalize this cautious operator assumption in the following section, since it aligns well with the observed one-sided fuzziness in the $I_{Y}$ dimension and the strict compliance to the distance rule $I_{D}$.
Figure 5: Real data plot w.r.t. the score components $X_{D}$ and $X_{Y}$. The decision boundary $T$ is dashed. The actual treatment assignment (red and blue) follows an unobserved rule $D$, rendering the MRD fuzzy.
A policymaker or manager might be interested in validating the decision policy $T$ in this application setting. We are using the subset identification from Section 3, to estimate effects at the cutoff. By evaluating these effects, we draw conclusions for improvements of the treatment assignment $T$.
4.2 Modeling assumptions and implications
In this section, we derive modeling assumptions on the decision rule $D$ for the application setting combining the insights from the observed data and the theory. Particularly, we assume that $D$ has an information advantage over the initial decision rule $T=I_{D}\land I_{Y}$. We make the assumption that this information advantage is due to the availability of more detailed information on the improvement score. Particularly, We assume that the initial treatment assignment $T$ is only due to the improvement estimates of every $m$-th item in the production lot resulting in the score $X_{Y}$ whereas the final decision-maker has access to the overall estimated yield improvement $X_{E}=X_{Y}+X_{R}$ including every item where $X_{R}$ is the yield improvement of items not taken into account by $X_{Y}$. See Algorithm 1 for details.
Data: seed, lot-size $n$, cutoff $c$, measurement steps $m$, yield criteria $\mathcal{Y}$, distance criteria $\mathcal{D}$, operator specific policy $D_{O}$
Result: lot $L$, scores $X=(X_{D},X_{Y})$, assigned treatment $T$, actual treatment $D$, outcome $Y$
$L\leftarrow(C_{1},\ldots,C_{n})$ generate a random production lot of $n$ items;
$X_{D}\leftarrow\mathcal{D}(L)$ calculate the distance to the target;
$\hat{L}_{A}\leftarrow$ carry out an optimal rework step on $L$;
$X_{Y}\leftarrow$ calculate improvement $\mathcal{Y}(\hat{L}_{A})-\mathcal{Y}(L)$ on every $m$-th item;
$T\leftarrow\operatorname{\mathbbm{1}}[X_{D}>c_{D}]\land\operatorname{\mathbbm{1}}[X_{Y}>c_{Y}]$;
$X_{E}\leftarrow$ calculate improvement $\mathcal{Y}(\hat{L}_{A})-\mathcal{Y}(L)$ on every item;
$D\leftarrow\operatorname{\mathbbm{1}}[X_{D}>c_{D}]\land D_{O}(X_{Y},\,X_{E})$;
$L_{A}\leftarrow$ carry out a realistic (noisy) rework step on $L$;
if D then
 $Y\leftarrow\mathcal{Y}(L_{A})$
else
 $Y\leftarrow\mathcal{Y}(L)$
end if
Algorithm 1 DGP for a knowledgeable operator
The implementation of this algorithm allows us to benchmark different operator assumptions and compare it to the real-world case. In semiconductor manufacturing the term operator refers to the human decision-maker carrying out $D$. Figure 6 displays exemplary data generated due to this DGP. While the knowledgeable operator has in principle access to a better improvement estimate, there are still several ways how she might apply this information. In the following we discuss several special cases of the knowledgeable operator assumption, allowing us (in theory) to identify the unit categories.
Figure 6: Comparison of data with cautious (left) and acknowledging (right) operator.
The acknowledging operator
In this scenario the operator $D$ accepts whatever treatment assignment $T$ has been made, despite better knowledge. Thus, the final treatment assignment is identical to the decision rule $D=T=I_{D}\land I_{Y}$ Which amounts to the operator specific policy in Algorithm 1 to $D_{O}=I_{Y}$. Consequently, one has $\operatorname{ComP}(T,D)=\mathcal{I}$ as well as
$\operatorname{ComP}(G,D)=\{i\,|\,X_{Y,i}>0\}\,\mbox{ and }\,\operatorname{Nt}(G,D)=\{i\,|\,X_{Y,i}\leq 0\}$
for $G=I_{D}$ using 7 or the Example 3.4).
The cautious operator
Now suppose that the operator is very considerate if it comes to an accidental degradation of the production lot due to the optional rework step. That is, whenever $T$ suggests that a rework step should be carried out, the operator may object it. This objection can be motivated due to better knowledge or further constraints (e.g. time) that the intended treatment rule $T$ does not capture. If $T$ does not suggest a rework treatment we make the assumption that the operator accepts this decision (e.g. due to the fear of possible degradation). We formalize the inclination of this negative overwrite into an additional score variable $X_{op}$ and model $D$ according to this assumption:
$D=(T\land I_{op})\lor(\overline{T}\land 0)=T\land I_{op}$
In particular we set $X_{op}=X_{E}$ and thus $D_{O}(X\,|\,c)\coloneqq I_{X_{Y}}\land I_{X_{E}}$. Thus, we have
$\operatorname{ComP}(T,D)=\{i\,|\,X_{E,i}>0\}\,\mbox{ and }\,\operatorname{Nt}(T,D)=\{i\,|\,X_{E,i}\leq 0\}$
as well as
$\operatorname{ComP}(G,D)=\{i\,|\,X_{E,i}>0,\,X_{Y}>0\}\,\mbox{ and }\,\operatorname{Nt}(G,D)=\{i\,|\,X_{E,i}\leq 0\}\cup\{i\,|\,X_{Y,i}\leq 0\}$
for the subrule $G\coloneqq I_{D}$ of $T$ again using Proposition 7 or the Example 3.4. Employing Theorem 2 one can estimate the sub-set complier effect of $G$ excluding the nevertaker $\Omega\coloneqq\{i\,|\,X_{Y,i}\leq 0\}$ instead of the complier effect of $G$. Since we assume that $X_{E}$ is only known by the operator the nevertaker due to $X_{E,i}\leq 0$ not excluded. Since the condition $X_{Y,i}\leq 0$ is global, the required continuity and stability assumptions are likely to be satisfied in practical applications.
The reasonable operator
Finally, imagine a operator that always uses his information advantage. That is, the reasonable operator overwrites the intended $T$ treatment using the full information $X_{E}$ instead of $X_{Y}$:
$D_{O}(X\,|\,c)=I_{E}(c)$
That means, the final decision rule $D$ can be related to $T$ by observing that:
$D(X\,|\,c)=I_{D}(c)\land\operatorname{\mathbbm{1}}[X_{Y}+X_{R}>c_{Y}]$
Thus $i\in\operatorname{ComP}(T,\,D)$ if $X_{R,i}=0$, which means that there is no overall improvement or degradation on the left out items. Now suppose that $X_{R,i}\neq 0$. Then we can choose $c_{Y}=2\max(X_{Y,i},X_{R,i})$, and $c_{D}<X_{D,i}$ to obtain $D(X_{i}\,|\,c)=T(X_{i}\,|\,c)=0$. But there exist also $c\in\operatorname{supp}(T)=\mathbb{R}^{2}$ such that $D(X_{i}\,|\,c)\neq T(X_{i}\,|\,c)$. For this set $c_{D}<X_{D,i}$. If $X_{R}\geq 0$ choose $c_{Y}=X_{Y}$ then $X_{Y}+X_{R}>c_{Y}$ and thus $T(X_{i}\,|\,c)=0$ and $D(X_{i}\,|\,c)=1$. Otherwise, choose $c=X_{Y}+X_{R}$ then $c<X_{Y}$ and thus $T(X_{i}\,|\,c)=1$ and $D(X_{i}\,|\,c)=0$. This shows that
$\operatorname{ComP}(T,\,D)=\{i\,|\,X_{R,i}=0\}$
and that every other item falls into the indecisive category. There are several shortcomings with this model. First, it does not reflect the observed data, since according to this model there should be treated items that did not get the initial treatment assignment. Second, the identification result relies on the absence of indecisive items. In the application small deviations of improvement score, that is small $X_{R}$, would also be considered compliant to the assignment $T$, requiring a local definition of the categories.
While still interesting for further theoretical investigations and as an example for the indecisive case we focus on the edge-cases “the acknowledging operator” and -“the cautious operator” as visible in Figure 6 in our investigation.
5 Simulation Study
In this section, we provide numerical results from the semi-synthetic data-generating process derived in the previous section. We aim at estimating the causal effect of the rework decision on the final yield in a neighborhood of the decision boundary. For this, we will use different estimators at both decision thresholds separately and make use of the identification theorems in Section 3.
5.1 Set-Up
For this, we generate semi-synthetic data following Algorithm 1. The python implementation of this process as well as all estimators is publicly available. The process is calibrated to match real data characteristics. We draw n=10,000n=10{,}000 observations and repeat each experiment $r=250$ times.
We evaluate the cut-offs $c_{D}$ and $c_{Y}$ separately, estimating the complier effect of $G\in\{I_{Y},\,I_{D}\}$ as identified in Theorem 1, the subset complier effect of $G$ given its counterpart and intent-to-treat estimates with and without the subset conditioning. The complier effects are estimated under a fuzzy design; the intent-to-treat effects under a sharp design.
The oracle values are estimated using local linear kernel regression on the differences in true potential outcomes. The covariates consist of statistics describing the quality of individual items. We use different estimators including covariate adjusted estimators estimating the complier and subset complier effect
Table 1 provides an overview of the estimators we use.
Method Adjustment
No Adjustment -
Conventional Adjustment Linear
RDFlex Lasso Local Penalized Linear Estimators
RDFlex Global Lasso Global Penalized Linear Estimators
RDFlex Boosting Gradient Boosting Estimators
RDFlex Stacking Stacked Combination of Linear and Non-Linear Estimators
Table 1: Considered methods and estimators in this section.
5.2 Estimators
The basic RDD estimator runs separated local linear regressions on each side of the cutoff:
$\hat{\tau}_{\text{base}}(h)=\sum_{i=1}^{n}w_{i}(h)Y_{i},$ (9)
where the $w_{i}(h)$ are local linear regression weights that depend on the data through the realizations of the running variable only and $h>0$ is a bandwidth.
Under standard conditions, which include that the running variable is continuously distributed, and that the bandwidth $h$ tends to zero at an appropriate rate, the estimator $\hat{\tau}_{\text{base}}(h)$ is approximately normally distributed in large samples, with bias of order $h^{2}$ and variance of order $(nh)^{-1}$.
The conventionally adjusted estimator includes covariates linearly in the local regressions. We also use modern RDD estimators with flexible covariate adjustment based on potentially nonlinear adjustment functions $\eta$. The estimator takes the following form:
$\widehat{\tau}_{\text{RDFlex}}(h;\eta)=\sum w_{i}(h)M_{i}(\eta),\quad M_{i}(\eta)=Y_{i}-\eta(Z_{i}).$ (10)
$\eta$ is the influence of $Z$ on the outcome $Y$ and is estimated by ML methods.
5.3 Cautious Operator
Figure 7 depicts a side-by-side the estimators for the effect at $c_{D}$ and $c_{Y}$ for the case of the cautious operator. The intent-to-treat oracles are closer to zero than the complier effects due to the inclusion of individuals who are nevertakers with respect to each cutoff rule. As visible in 7(a), for $I_{D}$ we estimate an overall negative effect, although it is not significant at $95\%$-level. The subset effect for the fuzzy case exhibits a smaller bias as the percentage of nevertakers in the estimation sample is smaller. The estimated effect remains unchanged as only nevertakers but no compliers were removed. For the intent-to-treat estimator, a higher share of compliers in the subsample increases the estimated effect of treatment rule $G$. The subset estimators have a comparable estimated variance (see Table 2), with the coverage overall appearing slightly more credible. Generally, the covariate adjustment reduces the standard error in the estimation, especially for the sharp estimators. The estimates for $I_{Y}$ are small and positive. The fuzzy estimator on the full data has a high standard error, which increases further with ML adjustment. This may be due to a small jump in treatment probability in the full data, destabilizing the ML estimate. The subset estimator along this axis removes much more observations within the bandwidth of the estimation thus decreasing the variance. Additional results can be found in Appendix B.
(a) Results for fuzzy and sharp estimators at the cutoff $c_{D}$ both on the full sample and on the subset.
(b) Results for fuzzy and sharp estimators at the cutoff $c_{Y}$ both on the full sample and on the subset.
Figure 7: Comparison of median estimates with median 95% confidence interval at $c_{D}$ and $c_{Y}$ on the simulated data with “cautious” operator assumption. Different colors depict complier effect (fuzzy) as well as intent-to-treat effect (sharp). The subset estimator excludes nevertaker according to Theorem 2.
Mean Bias s.e. Coverage RMSE left Log loss left RMSE right Log loss right
setting method
Fuzzy RDD Conventional Covs -0.0025 0.0111 0.9912
RDD Without Covs -0.0027 0.0109 0.9735
RDFlex Boosting -0.0025 0.0095 0.9912 0.0245 0.0532 0.0923 0.4996
RDFlex Global Lasso -0.0028 0.0097 0.9823 0.0399 0.0987 0.0897 0.6070
RDFlex Lasso -0.0030 0.0092 0.9823 0.0242 0.0787 0.0897 0.5645
RDFlex Stacking -0.0029 0.0092 0.9823 0.0243 0.0169 0.0898 0.4938
Fuzzy Subset RDD Conventional Covs -0.0001 0.0117 0.9609
RDD Without Covs -0.0003 0.0118 0.9478
RDFlex Boosting -0.0009 0.0116 0.9652 0.0305 0.0741 0.1050 0.3786
RDFlex Global Lasso -0.0004 0.0111 0.9522 0.0259 0.1961 0.1012 0.3665
RDFlex Lasso -0.0002 0.0110 0.9435 0.0249 0.4230 0.1010 0.5140
RDFlex Stacking -0.0004 0.0113 0.9522 0.0258 0.0329 0.1011 0.3580
Sharp RDD Conventional Covs -0.0040 0.0063 0.9120
RDD Without Covs -0.0040 0.0063 0.9240
RDFlex Boosting -0.0040 0.0056 0.9080 0.0247 0.0918
RDFlex Global Lasso -0.0040 0.0054 0.9240 0.0401 0.0888
RDFlex Lasso -0.0040 0.0055 0.9200 0.0242 0.0886
RDFlex Stacking -0.0040 0.0055 0.9240 0.0243 0.0888
Sharp Subset RDD Conventional Covs -0.0014 0.0094 0.9680
RDD Without Covs -0.0014 0.0095 0.9600
RDFlex Boosting -0.0016 0.0085 0.9760 0.0280 0.1045
RDFlex Global Lasso -0.0014 0.0079 0.9640 0.0253 0.1006
RDFlex Lasso -0.0014 0.0079 0.9560 0.0246 0.1004
RDFlex Stacking -0.0015 0.0080 0.9600 0.0256 0.1006
Table 2: Mean Bias and standard error, coverage and first stage prediction quality of different estimators in estimation at $c_{D}$.
Two-dimensional Estimators
It is worth mentioning some results of two-dimensional estimators as suggested in previous literature. We compare the “binding-the-score” approach, that normalizes and combines $X_{D}$ and $X_{Y}$ into a singular score component and the “Euclidean distance” approach, that computes the shortest distance of each observation to a predetermined point (in our case a point on $c_{D}$.
Figure 8: Median estimates with median 95% confidence interval Comparison of median estimates with median 95% confidence interval Estimators that use two dimensions at once.
The results in Figure 8 show oracles very close to zero. This indicates that when collapsing both dimensions of the decision rule into one it is no longer possible to have individual insights into the rules. Interestingly, the estimator that uses the euclidean distance to a boundary point exhibits a significant bias across all methods. In our specific setting, there is no straightforward interpretation of the RDD effect estimate at a two-dimensional threshold, which contrasts with previous work where score components are more directly comparable, such as test scores or geographical distances.
5.4 Acknowledging Operator
We compare the results from the previous section to the “acknowledging” operator, who accepts the the decision of the system without using his additional domain knowledge. Thus, in this setting the data follows a sharp two-dimensional MRD. The comparison of the effects at $c_{Y}$ and $c_{D}$ can be seen in Figure 9
(a) Results for fuzzy and sharp estimators at the cutoff $c_{D}$ both on the full sample and on the subset.
(b) Results for fuzzy and sharp estimators at the cutoff $c_{Y}$ both on the full sample and on the subset.
Figure 9: Comparison of median estimates with median 95% confidence interval at $c_{D}$ and $c_{Y}$ on the simulated data with “acknowledging” operator assumption. Different colors depict complier effect (fuzzy) as well as intent-to-treat effect (sharp). The subset estimator excludes nevertaker according to Theorem 2
In this scenario, the effects are smaller for $c_{D}$, as lots that the cautious operator did not rework due to his domain knowledge about possible degradation, are reworked. This leads to an overall larger negative effect and suggests moving the threshold even further than in the scenario above. The sharp subset estimator estimates the identical effect as the fuzzy estimator on the full data, as subsetting the data according to $I_{Y}$ leads to a sharp design with all individuals complying to threshold movements. The subset estimators estimate the effect overall with a slightly smaller bias and standard error. The sharp estimator on the full sample, which corresponds to an intent-to-treat estimator, estimates an effect closer to zero, as it includes nevertakers.
At the cutoff $c_{Y}$, the change of behavior of the operator does not change the effect estimates much. This can be explained that less individuals in the neighborhood of the cutoff are effected by this.
6 Empirical Results
This section presents results from estimation on the real data with n=9,103n=9{,}103 observations from the production system. Additional to the covariates used above, we include the shopfloor workload. Our primary interest in the real system is evaluating $I_{D}$.
As shown in Table 3, we estimate a positive effect. The subset estimates have a lower variance, as we remove uncertainty due to nevertakers according to $I_{Y}$. In case of the intent-to-treat estimators, the addition of ML adjustment improves the variance of the estimation. For the fuzzy estimators, there is no clear improvement.
Coef s.e. CI 2.5% CI 97.5% RMSE left Log loss left RMSE right Log loss right % s.e. change
setting method
Fuzzy RDD Conventional Covs 0.1031 0.0508 0.0035 0.2028 22.2532
RDD Without Covs 0.0387 0.0416 -0.0428 0.1202 0.0000
RDFlex Boosting 0.0741 0.0524 -0.0400 0.1882 0.1432 0.0222 0.1594 0.6264 25.9991
RDFlex Global Lasso 0.0519 0.0461 -0.0484 0.1522 0.1597 0.0108 0.1510 0.6653 10.8590
RDFlex Local Lasso 0.0999 0.0470 -0.0004 0.2003 0.1452 0.0141 0.1523 0.6142 12.9249
RDFlex Stacking 0.1097 0.0500 0.0023 0.2171 0.1429 0.0188 0.1570 0.6322 20.1333
Fuzzy Subset RDD Conventional Covs 0.0630 0.0426 -0.0205 0.1466 36.5324
RDD Without Covs 0.0369 0.0312 -0.0243 0.0981 0.0000
RDFlex Boosting 0.0581 0.0332 -0.0163 0.1325 0.1423 0.0200 0.1562 0.5158 6.4133
RDFlex Global Lasso 0.0270 0.0375 -0.0537 0.1076 0.1548 0.0113 0.1520 0.5731 20.2628
RDFlex Local Lasso 0.0713 0.0304 0.0031 0.1395 0.1410 0.0209 0.1507 0.5424 -2.5667
RDFlex Stacking 0.0768 0.0319 0.0048 0.1487 0.1394 0.0121 0.1530 0.5018 2.1538
Sharp RDD Conventional Covs 0.0393 0.0208 -0.0015 0.0802 2.2467
RDD Without Covs 0.0165 0.0204 -0.0235 0.0564 0.0000
RDFlex Boosting 0.0311 0.0176 -0.0093 0.0715 0.1417 0.1575 -13.5386
RDFlex Global Lasso 0.0200 0.0171 -0.0193 0.0593 0.1545 0.1512 -16.3636
RDFlex Local Lasso 0.0357 0.0167 -0.0026 0.0740 0.1425 0.1513 -17.9717
RDFlex Stacking 0.0367 0.0168 -0.0018 0.0753 0.1408 0.1530 -17.5203
Sharp Subset RDD Conventional Covs 0.0362 0.0205 -0.0040 0.0764 -2.3944
RDD Without Covs 0.0098 0.0210 -0.0314 0.0509 0.0000
RDFlex Boosting 0.0222 0.0196 -0.0225 0.0669 0.1424 0.1590 -6.6630
RDFlex Global Lasso 0.0102 0.0179 -0.0306 0.0509 0.1533 0.1518 -14.7987
RDFlex Local Lasso 0.0364 0.0184 -0.0031 0.0758 0.1408 0.1525 -12.2509
RDFlex Stacking 0.0307 0.0181 -0.0093 0.0708 0.1395 0.1540 -14.0691
Table 3: Coefficient, standard error, and quality of fit for the different estimators for $c_{D}$ in the real data.
The comparison in Figure 10 provides insight into the differing effect signs between semi-synthetic and real data. While the semi-synthetic process (left) appears to have a rework threshold closer than the optimal distance, the real data (right) suggests a threshold that is further away, resulting in a significant outcome jump at the current threshold.
(a) RDD plot for the semi-synthetic process.
(b) RDD plot for the real production data.
Figure 10: Comparison of local linear regressions around the cutoff for the semi-synthetic and real data. The opposite sign of the estimated effect can be explained by the calibration of the semi-synthetic process.
Validation Check.
We validate our results from the application with a pseudo-cutoff test (see Cattaneo et al. (2019)). We compare the estimated intent to treat effect of the real data at $c_{D}$ with a fictional lower and a higher cutoff.
In Table 4 it is visible that both pseudo cutoffs show effects close to zero and with large confidence band, while at the real cutoff the effect is almost significant on $95\%$ confidence level.
Cutoff Coef 2.5 % CI 97.5 % CI
lower pseudo 0.002919 -0.014796 0.020634
real 0.039332 -0.001524 0.080188
higher pseudo 0.041137 -0.015333 0.097606
Table 4: Coefficient and confidence interval of the sharp real data estimation at pseudo cutoffs concerning $c_{D}$.
7 Conclusion
With this paper we present progress towards better applicability of the regression discontinuity design in Economics and Management. The multi-score RDD framework combined with modern ML adjustment yields great opportunities to evaluate complex decision settings and policies. Particularly, in settings with strict decision boundaries in place, policy makers experience little to no overlap, limiting the possibility of “what happens if”-reasoning.
Our theoretical insights are based on three intuitive observations. First, the analysis of more general assignment rules requires lifting the concept of “binary compliance” to the higher-level concept of “compliance with an assignment mechanism”. That is, in theory all possible inputs of the mechanism need to be considered when assessing compliance. Secondly, and most evident by the first observation, the concept of compliance is only relative to the assessed rules and thus changes of the assignment mechanism can alter the perceived unit behavior. A fact that should be reflected in the theory as knowledge about parts of the assignment mechanism and/or the final decision-rule allows identifying noncompliance. Lastly, in multiple dimensions the complier effect at the cutoff should not depend on the direction from which it is approached, a property that can be exploited by future estimation methods. Formalizing these observations we generalized existing unit behavior types (complier, nevertaker, alwaystaker) for multi-dimensional cutoff-rules fully characterizing the possible behavior types with the introduction of indecisive items. We derived rules that show how the unit behavior changes when cutoff-rules are decomposed into sub-rules. Further we proved an identification result for the cutoff complier effect in MRD settings, and provided conditions for the stability of this result when subsets of alwaystaker and nevertaker are removed, giving rise to the sub-set complier effect at the cutoff.
We validated our theory both with simulation studies using a semi-synthetic DGP and in a real-world application setting from manufacturing, showing that that subrule-specific estimation provides insights on how to improve existing decision-making. In particular the results confirm that removing identifiable nevertaker and/or alwaystaker allow for more precise effect estimation, under conditions provided by the theory. Combined with flexible ML-based adjustments variance can be further reduced without introducing additional bias. We advise the practitioner to estimate both the complier and the sub-set complier effect at the cutoff whenever nevertaker and/or alwaystaker are identifiable in application, and retrospectively validate the assumptions. Importantly, the empirical findings underline the asymmetry in the treatment effects across different cutoff dimensions, highlighting that effect of the whole assignment rule is in general not equal to the sub-rule effects, as is evident by the theory. Nevertheless, a nuanced evaluation of each score component, can provide valuable insights.
We expanded the existing MRD theory providing means for the analysis of complex cutoff-rules, and their corresponding sub-rules. For this we used the language of well-defined unit behavior types. While we bridged the gap between econometric RDD methods and industrial process control, our results remain broadly applicable in complex decision settings. Future work could explore the relation of main- and sub-rules more closely, expand on the stability of unit types, or extend the multi-dimensional RDD setting to hierarchical decision structures across multiple stages. Additionally, developing estimators tailored to the identification theorems are promising future avenues.
Appendix A Proofs
A.1 Unit categorization
See 1 Before proofing the above statement let us make some simple observations. Let
$S(T\,|\,X)\coloneqq\{k\in\mathbb{N}_{\leq K}\,|\,\lambda\in\mathbb{R}\,:\,T_{i}(X\,|\,0)\neq T(X\,|\,0+\lambda e_{k})\}$
denote the set of local change directions of $X$. Then one has the following properties:
  1. 1.
    S(T|X)S(T\,|\,X) depends only upon the quadrant of $X$ since within a quadrant the $I_{k}$‘s and $\overline{I}_{k}$‘s are constant.
  2. 2.
    $S(T)=\bigcup_{X\in\mathbb{R}^{K}}S(T\,|\,X)$
    Since $T(0\,|\,c+\lambda e_{k})\neq T(0\,|\,c)$ is equivalent to $T(-c\,|\,0+\lambda e_{k})\neq T(-c\,|\,0)$.
  3. 3.
    S(T|X)S(T\,|\,X) can be empty.
    For example let $T=(I_{1}\land I_{2})\lor I_{3}$ and $X\in\mathbb{R}^{3}_{>0}$. Then $S(T\,|\,X)=\emptyset$, since $T$ is true as long as there are at least two indicators that are true. Changing the cutoff in only one direction $k$ does not affect $T(0+\lambda e_{k})$. This example can be extended such that changes in multiple directions do not affect $T$.
Proof.
One direction is clearly trivial. For the other suppose $T$ is not constant. Then one has $0=T(X\,|\,0)\neq T(\hat{X}\,|\,0)=1$ for appropriate $X,\,\hat{X}\in\mathbb{R}^{K}$. Define the sequence
$X^{k}\coloneqq X+\sum_{j=1}^{k}\left\langle\hat{X}-X,\,e_{j}\right\rangle e_{j}\quad 0\leq k\leq K$
for $0\leq k\leq K$. Then $X^{K}=\hat{X},\,X^{0}=X$ and $X^{k}-X^{k-1}=\lambda_{k}e_{k}$ with $\lambda_{k}\coloneqq\left\langle\hat{X}-X,\,e_{k}\right\rangle$. We assume that $S(T\,|\,X^{k})=\emptyset$ for all $k$. Then
$T(X^{k-1}\,|\,0)=T(X^{k-1}\,|\,0-\lambda_{k}e_{k})=T(X^{k}\,|\,0)$
and thus per induction 0=T(X| 0)=T(X^| 0)=10=T(X\,|\,0)=T($\hat${X}\,|\,0)=1. Which contradicts the assumption. Thus, there exists some $k$ with $S(T\,|\,X^{k})\neq\emptyset$. Which shows that $S(T)\neq\emptyset$. ∎
As a general assumption for the next statements we require that $T$ is not trivial, that is $\operatorname{supp}(T)\neq\{0\}$. See 1
Proof.
Using equations 3 one can see that $X+\lambda Y\in N^{T}$ for $X,Y\in N^{T}$ and $\lambda\in\mathbb{R}_{>0}$. Further T(X|c)=T(0|c+X)=T(X|c+X)=T(XX|c)=T(0|c)T(-X\,|\,c)=T(0\,|\,c+X)=T(X\,|\,c+X)=T(X-X\,|\,c)=T(0\,|\,c) for $X\in N^{T}$ and $c\in\mathbb{R}^{K}$. Thus, $N^{T}$ is a linear subspace of $\mathbb{R}^{K}$. Let $k\in S(T)$ then there exists $\lambda\in\mathbb{R},\,c\in\mathbb{R}^{K}$ such that $T(0\,|\,c+\lambda e_{k})\neq T(0\,|\,c)$. This means that $\lambda e_{k}\notin N$. Since $N^{T}$ is a linear space this means that $e_{k}$ can also not be in $N$. Now let $k\notin S(T)$. Then $T(0\,|\,c)=T(0\,|\,c-e_{k})=T(e_{k}\,|\,c)$ for all $c\in\mathbb{R}^{K}$ and thus $e_{k}\in N^{T}$. This shows that $\operatorname{supp}(T)\cap N^{T}=\{0\}$. Let $X^{T}\coloneqq\sum_{k\in S(T)}\left\langle X,\,e_{k}\right\rangle e_{k}$ be the projection on to the subspace $\operatorname{supp}(T)$ and $X^{\perp T}\coloneqq X-X^{T}$. Suppose there exists some $c$ such that $T(X^{\perp T}\,|\,c)\neq T(0\,|\,c)$. Let
$c^{k}\coloneqq c-\hat{X}+\sum_{j=1}^{k}\lambda_{j}e_{j}\quad 0\leq k\leq K$
with $\lambda_{j}\coloneqq\left\langle X^{\perp T},\,e_{j}\right\rangle$. Then $c^{0}=c-X^{\perp T},\,c^{K}=c$ and $c^{k}=c^{k-1}+\lambda_{k}e_{k}$ holds. Further, one has
$T(0\,|\,c^{k})=T(0\,|\,c^{k-1}+\lambda_{k}e_{k})=T(0\,|\,c^{k-1})$
for all $1\leq k\leq K$. Otherwise, $k\subset S(T)$ and thus $\lambda_{k}=0$ by definition of $X^{\perp T}$. Using induction we derive $T(X^{\perp T}\,|\,c)=T(0\,|\,c^{0})=T(0\,|\,c^{K})$ which contradicts the assumption. For uniqueness suppose there exists another decomposition $X=Z^{T}+Z^{\perp T}$ with the above properties. Then $Z^{T}-X^{T}=Z^{\perp T}-X^{\perp T}\in\operatorname{supp}(T)\cap N^{T}=\{0\}$. ∎
See 2
Proof.
[…]
See 3
Proof.
It is easy to see that $\operatorname{At}(T,D)\cap\operatorname{Nt}(T,D)=\emptyset$ and $\operatorname{ComP}(T,D)\cap\operatorname{DeF}(T,D)=\emptyset$. For $i\in C\coloneqq\operatorname{Nt}(T,D)\cup\operatorname{At}(T,D)$ it follows that $D(X_{i}^{\perp T}-c)$ is constant for all $c\in\operatorname{supp}(T)$. Suppose that $i\in C\cap\operatorname{ComP}(T,D)$. The latter would mean, that $T(0\,|\,c)=D(X_{i}^{\perp T}-c)$ for all $c\in\operatorname{supp}(T)$. This contradicts $\operatorname{supp}(T)\neq\{0\}$. Now suppose that $i\in C\cap\operatorname{DeF}(T,D)$. The latter would mean that $T(0\,|\,c)\neq D(X_{i}^{\perp T}-c)$ for all $c\in\operatorname{supp}(T)$, which implies that $T$ is constant on $\operatorname{supp}(T)$. This again contradicts $\operatorname{supp}(T)\neq\{0\}$. ∎
From now on whenever we assume that $D$ is a cutoff rule we suppose that $D$ does not depend on $I_{k,i}(c)$ with $c\neq 0$, that is $T$ and $D$ are synchronous regarding their cutoffs. See 4
Proof.
To show the first implication we suppose that $\dim(\operatorname{supp}(T))>1$ and construct a $D$ such that $\operatorname{Ind}(T,\,D)\neq\emptyset$. Since $\dim(\operatorname{supp}(T))>1$ here exist $l,k\in S(T)$ with $l\neq k$. Define
$D=\left(T\land I_{k}\land I_{l}\right)\lor\left(\overline{T}\land\overline{I_{k}}\land I_{l}\right)\lor\left(I_{k}\land\overline{I_{l}}\right)$
and let $X_{i}$ be such that $X_{k,i}=0=X_{l,i}$. Then $D(X_{i}\,|\,0)=0$ and $D(X_{i}\,|\,e_{k})=1$. This shows that $i\notin\operatorname{At}(T,\,D)\cup\operatorname{Nt}(T,\,D)$ since $0,\,e_{k}\in\operatorname{supp}(T)$. Further
$D(X_{i}\,|\,e_{k}+e_{l})=T(X_{i}\,|\,e_{k}+e_{l})$
and
$D(X_{i}\,|\,e_{l})=\overline{T}(X_{i}\,|\,\hat{c})\neq T(X_{i}\,|\,e_{l}).$
Thus, $i\in\operatorname{Ind}(T,\,D)$ since $e_{l},\,e_{k}\in\operatorname{supp}(T)$.
[…]
A.2 Examples
Example (OR-Rules)
Let $D\coloneqq\bigvee_{j=1}^{K}I_{j}$ and $T\coloneqq\bigvee_{j=1}^{k}I_{j}$ for $k\in\mathbb{N}_{<K}$. Then $\operatorname{supp}(T)=\mathbb{R}^{k}\times\{0\}^{K-k}$. The potential outcomes reduce to $Y_{i}=Y_{i}(0,0)(1-D_{i})+Y_{i}(0,1)D_{i}(1-T_{i})+Y(1,1)T_{i}$. Further we have the following unit categorizations:
  1. 1.
    $\operatorname{ComP}(T,\,D)=\{i\,|\,X_{i}\in\mathbb{R}^{k}\times\mathbb{R}^{K-k}_{\leq 0}\}$
    Note that in this case the additional or-conditions are zero, reducing the rule $D_{i}$ to $T_{i}$.
  2. 2.
    $\operatorname{At}(T,\,D)=\{i\,|\,\exists k<j\leq K\,:\,X_{j,i}>0\}=\{i\,|\,X_{i}\in\mathbb{R}^{k}\times(\mathbb{R}^{K-k}\setminus\mathbb{R}^{K-k}_{\leq 0})\}$
    As long as there is any additional or condition that is always true, cutoff changes in $\operatorname{supp}(T)$ do not affect $D$.
  3. 3.
    $\operatorname{Nt}(T,\,D)=\emptyset$
    Choose $c\in\operatorname{supp}(T)$ such that $X_{i,j}>c$ for $0\leq j\leq k$, then $D_{i}=1$.
  4. 4.
    $\operatorname{DeF}(T,\,D)=\emptyset$
    Note that $T_{i}(c)=1$ implies $D_{i}(c)=1$ for all $c\in\operatorname{supp}(T)$.
Example (XOR dominant rule)
Let $D\coloneqq(I_{1}\lor I_{2})\land(\overline{I}_{1}\lor\overline{I}_{2})$ and $T\coloneqq I_{1}$. Then $\operatorname{supp}(T)=\mathbb{R}\times\{0\}$. The potential outcomes formulation does not simplify much:
$Y_{i}=Y_{i}(0,0)(1-T_{i})(1-I_{2,i})+Y_{i}(1,0)T_{i}I_{2,i}+Y_{i}(0,1)(1-T_{i})I_{2,i}+Y(1,1)T_{i}(1-I_{2,i})$
Note if $X_{2,i}>0$, one has $D_{i}(c)=\overline{T}_{i}(c)$ and otherwise $D_{i}(c)=T_{i}(c)$ for all $c\in\operatorname{supp}(T)$. Thus:
  1. 1.
    $\operatorname{ComP}(T,\,D)=\{i\,|\,X_{2,i}\leq 0\}$
  2. 2.
    $\operatorname{At}(T,\,D)=\emptyset$
  3. 3.
    $\operatorname{Nt}(T,\,D)=\emptyset$
  4. 4.
    $\operatorname{DeF}(T,\,D)=\{i\,|\,X_{2,i}>0\}$
Nevertaker, Complier, Alwaystaker
We now investigate potential setup for cutoff rules that admit complier, alwaystaker and nevertaker. In accordance with the empirical part let $T\coloneqq I_{1}\square I_{2}$ for $\square\in\{\lor,\land\}$ and $G=I_{k}$ for $k=1,2$ and further let $D$ be a cutoff rule. Thus, only the number of additional variables in $D$ limit the possible unit categories. That is, with one additional cutoff variable $I_{3}$ one could realize each tuple of the different item types (e.g. defier and nevertaker). A common assumption in causal literature is that there are no defying or indecisive units. Thus, to have a setting with each of the remaining three categories one would require one more variable $I_{4}$. This leaves a single choice open (nevertaker, alwaystaker or complier) for the remaining region in $\mathbb{R}^{4}$. As in case of the cautious operator we decide to opt for compliance. Thus, we arrive at
$D=(T\land I_{3})\lor(\overline{I}_{3}\land I_{4})$
for a model of the actual treatment decision with:
$\operatorname{ComP}(T,D)$ $=\{i\,|\,X_{3,i}>0\}$
$\operatorname{At}(T,D)$ $=\{i\,|\,X_{3,i}\leq 0,X_{4,i}>0\}$
$\operatorname{Nt}(T,D)$ $=\{i\,|\,X_{3,i}\leq 0,X_{4,i}\leq 0\}$
$\operatorname{DeF}(T,D)$ $=\emptyset$
For the case $\square=\land$ and $G\coloneqq I_{1}$ one has
$\operatorname{ComP}(G,D)$ $=\{i\,|\,X_{3,i}>0,X_{2,i}>0\}$
$\operatorname{At}(G,D)$ $=\{i\,|\,X_{3,i}\leq 0,X_{4,i}>0\}$
$\operatorname{Nt}(G,D)$ $=\{i\,|\,X_{3,i}\leq 0,X_{4,i}\leq 0\}\cup\{i\,|\,X_{2,i}\leq 0,X_{4,i}\leq 0\}\cup\{i\,|\,X_{2,i}\leq 0,X_{3,i}>0\}$
$\operatorname{DeF}(G,D)$ $=\emptyset$
That is $X_{4}$ could be ignored for estimating the complier effect of $G$ due to our the specific choice, whereas $X_{3}$ and $X_{2}$ are the switches required to identify $\operatorname{ComP}(G,D)$ completely. Again $X_{2}$ might be known from $T$, which helps to reduce the observational dataset accordingly.
A.3 Inheritance
The following lemma is the basis for duality statements involving the unit categories.
Lemma 2.
The decomposition in Proposition 1 is stable under negation, that is
[…]
In particular $\operatorname{supp}(T)=\operatorname{supp}(\overline{T})$ and $N^{T}=N^{\overline{T}}$.
Proof.
If $T(0\,|\,c)\neq T(0\,|\,c+\lambda e_{k})$ holds, the same is true for $\overline{T}$ and vice versa. This shows that $S(T)=S(\overline{T})$ and $\operatorname{supp}(T)=\operatorname{supp}(\overline{T})$. Now let $X\in\mathbb{R}^{K}$ such that T(X|c)=T(X| 0)T(X\,|\,c)=T(X\,|\,0) for all $c\in\mathbb{R}^{K}$. Then the same equality holds for $\overline{T}$ and vice versa. Thus, $N^{T}=N^{\overline{T}}$. Using that the decomposition in Proposition 1 is a direct sum of linear spaces we get
$X_{i}^{T}-X_{i}^{\overline{T}}=X_{i}^{\perp\overline{T}}-X_{i}^{\perp T}\in\operatorname{supp}(T)\cap N^{T}=\{0\}$
which concludes the proof. ∎
See 5
Proof.
Using Lemma 2 we have $D(X_{i}^{\perp T}-c)=D(X_{i}^{\perp\overline{T}}-c)$ for $X_{i}\in\mathbb{R}^{K}$ and $c\in\operatorname{supp}(T)=\operatorname{supp}(\overline{T})$. The equalities in of Definition 3 are easily checked. ∎
See 6
Proof.
One has $(X_{i}^{\perp G})^{\perp T}=X_{i}^{\perp T}$. To see this note that $X_{i}=(X_{i}^{\perp G})^{T}+X_{i}^{G}+(X_{i}^{\perp G})^{\perp T}$ with $(X_{i}^{\perp G})^{T}+X_{i}^{G}\in\operatorname{supp}(T)$. Then the claim follows by the uniqueness of the decomposition according to Proposition 1 . With this one obtains
$D(X_{i}^{\perp G}-c^{G})=D\left((X_{i}^{\perp G})^{\perp T}+(X_{i}^{\perp G})^{T}-c^{G}\right)=D(X_{i}^{\perp T}-c)$
with $c=c^{G}-(X_{i}^{\perp G})^{T}\in\operatorname{supp}(T)$ which concludes the proof. ∎
Lemma 3.
Let G,HG,\,H be cutoff rules on $\mathbb{R}^{K}$ with $\operatorname{supp}(T)=\operatorname{supp}(G)\oplus\operatorname{supp}(H)$. Then $N^{G}=\operatorname{supp}(H)\oplus N^{T}$.
Proof.
Let $X\in\operatorname{supp}(H)\cap N^{T}=\{0\}$, then $X\in\operatorname{supp}(T)$ since $\operatorname{supp}(H)\subset\operatorname{supp}(T)$ and thus $X\in N^{T}\cap\operatorname{supp}(T)=\{0\}$. This shows that the sum on the right side is indeed a direct sum, that is $\operatorname{supp}(H)+N^{T}=\operatorname{supp}(H)\oplus N^{T}$. Employing Proposition 1 one has $N^{G}\oplus\operatorname{supp}(G)=\mathbb{R}^{K}$. Using this and $\operatorname{supp}(G)\subset\operatorname{supp}(T)$ it follows that $N^{T}\subset N^{G}$. Also note that $\operatorname{supp}(H)\cap\operatorname{supp}(G)=\{0\}$ and thus $\operatorname{supp}(H)\subset N^{G}$. This shows $\operatorname{supp}(H)\oplus N^{T}\subset N^{G}$ since $N^{G}$ is a linear space. Now let $X\in N^{G}$. With Proposition 1 one has $X=X^{\perp T}=X^{T}+X^{\perp T}$ with $X^{T}\in\operatorname{supp}(T)$ and $X^{\perp T}\in N^{T}\subset N^{G}$ and further $X^{T}=X^{G}+X^{H}$ with $X^{G}\in\operatorname{supp}(G)$ and $X^{H}\in\operatorname{supp}(H)\subset N^{G}$. Thus, $X^{G}=X-X^{H}-X^{\perp T}\in\operatorname{supp}(G)\cap N^{G}$ and with this $X=X^{H}+X^{\perp T}\in\operatorname{supp}(H)\oplus N^{T}$. ∎
See 7
Proof.
We only show the case $T=G\land H$ since the other follows in the same way. Let $i\in\operatorname{ComP}(G,T)$. This is equivalent to
$G(X_{i}\,|\,c^{G})=G(X_{i}\,|\,c^{G})\land H(X_{i}\,|\,c^{G})=G(X_{i}\,|\,c^{G})\land H(X_{i}^{H}\,|\,0)\quad\forall\,c^{G}\in\operatorname{supp}(G)$
where the last equality follows from $\operatorname{supp}(G)\subset N^{H}$ (Proposition 1). Since the support of $G$ is non-trivial this is equivalent to $H(X_{i}^{H}\,|\,0)=1$. Now let $i\in\operatorname{Nt}(G,T)$. This is equivalent to:
$0=G(X_{i}\,|\,c^{G})\land H(X_{i}^{H}\,|\,0)\quad\forall\,c^{G}\in\operatorname{supp}(G)$
Again since the support of $G$ is non-trivial this is equivalent to $H(X_{i}^{H}\,|\,0)=0$. ∎
See 8
Proof.
Let $T=G\land H$ and $X_{i}\in\operatorname{ComP}(T,D)$ then
$G(0\,|\,c^{G})\land H(0\,|\,c^{H})=G(0\,|\,c)\land H(0\,|\,c)=T(0\,|\,c)=D(X_{i}^{\perp T}-c)$
for all $c=c^{G}+c^{H}\in\operatorname{supp}(G)\oplus\operatorname{supp}(H)=\operatorname{supp}(T)$. The first equality follows from $c^{H}\in N^{G}$ and $c^{G}\in N^{H}$ using Lemma 3. The above also holds for $c=c^{G}-X^{H}_{i}$. Further $X_{i}^{\perp G}=X_{i}^{H}+X_{i}^{\perp T}$ according to Lemma 3 and with this
$G(0\,|\,c^{G})\land H(X_{i}^{H}\,|\,0)=D(X^{\perp G}_{i}-c^{G})$
for all $c^{G}\in\operatorname{supp}(G)$. If $H(X_{i}^{H}\,|\,0)=1$ then it follows that $i\in\operatorname{ComP}(G,D)$. Otherwise, $D(X_{i}^{\perp G}-c^{G})=0$ for all $c^{G}\in\operatorname{supp}(G)$. Since $\{0\}\neq\operatorname{supp}(G)$ one has $i\in\operatorname{Nt}(G,D)$. In summary, using Lemma 7 one derives
$\operatorname{ComP}(G,T)\cap\operatorname{ComP}(T,D)$ $\subset\operatorname{ComP}(G,D)$
$\operatorname{Nt}(G,T)\cap\operatorname{ComP}(T,D)$ $\subset\operatorname{Nt}(G,D)$
and with it:
$\operatorname{ComP}(T,D)\subset\operatorname{ComP}(G,D)\cup\operatorname{Nt}(G,D)$
Using the same argument one derives the analog result for $T=G\lor H$. ∎
Appendix B Additional Numerical Results
In this appendix we present additional results and supplementary material to the simulation results in Section 5.3.
B.1 Simulation for Cautious Operator
Cutoff $c_{Y}$
Mean Bias s.e. Coverage RMSE left Log loss left RMSE right Log loss right
setting method
Fuzzy RDD Conventional Covs -0.0158 0.0124 0.6867
RDD Without Covs -0.0189 0.0117 0.5200
RDFlex Boosting 0.2453 7.8523 0.9467 0.0268 0.0313 0.0741 0.4072
RDFlex Global Lasso 0.1593 9.1373 0.9800 0.0374 0.0132 0.0707 0.5320
RDFlex Lasso -13.0991 690.4369 0.9600 0.0275 0.0202 0.0707 0.4681
RDFlex Stacking -0.3166 16.4743 0.9667 0.0257 0.0104 0.0732 0.3574
Fuzzy Subset RDD Conventional Covs -0.0003 0.0085 0.9754
RDD Without Covs -0.0002 0.0085 0.9713
RDFlex Boosting -0.0002 0.0060 0.9713 0.0359 0.1205 0.0950 0.1473
RDFlex Global Lasso 0.0002 0.0057 0.9836 0.0336 0.1394 0.0936 0.1468
RDFlex Lasso 0.0002 0.0057 0.9795 0.0333 0.2688 0.0936 0.1580
RDFlex Stacking 0.0000 0.0058 0.9754 0.0337 0.0321 0.0938 0.1460
Sharp RDD Conventional Covs -0.0043 0.0073 0.8720
RDD Without Covs -0.0050 0.0074 0.8880
RDFlex Boosting -0.0040 0.0059 0.8640 0.0257 0.0749
RDFlex Global Lasso -0.0049 0.0075 0.8840 0.0353 0.0728
RDFlex Lasso -0.0045 0.0077 0.8800 0.0271 0.0726
RDFlex Stacking -0.0048 0.0070 0.8680 0.0255 0.0732
Sharp Subset RDD Conventional Covs -0.0005 0.0078 0.9800
RDD Without Covs -0.0004 0.0078 0.9720
RDFlex Boosting -0.0007 0.0057 0.9840 0.0361 0.0953
RDFlex Global Lasso -0.0003 0.0054 0.9800 0.0337 0.0940
RDFlex Lasso -0.0003 0.0054 0.9760 0.0333 0.0940
RDFlex Stacking -0.0004 0.0054 0.9720 0.0337 0.0939
Table 5: Mean Bias and standard error, coverage and first stage prediction quality of different estimators in estimation at $c_{Y}$.
Two-dimensional estimator
Mean Bias s.e. Coverage RMSE left Log loss left RMSE right Log loss right
setting method
2D Binding Fuzzy RDD Conventional Covs -0.0032 0.0092 0.5080
RDD Without Covs -0.0020 0.0092 0.5080
RDFlex Boosting -0.0023 0.0073 0.5160 0.0292 0.0658 0.1008 0.2655
RDFlex Global Lasso -0.0029 0.0078 0.5120 0.0416 0.1128 0.0993 0.3470
RDFlex Lasso -0.0022 0.0074 0.5160 0.0288 0.3225 0.0987 0.4954
RDFlex Stacking -0.0023 0.0073 0.5120 0.0281 0.0155 0.0988 0.2438
2D Binding Sharp RDD Conventional Covs -0.0021 0.0081 0.9600
RDD Without Covs -0.0008 0.0081 0.9480
RDFlex Boosting -0.0019 0.0069 0.9720 0.0298 0.1011
RDFlex Global Lasso -0.0018 0.0068 0.9520 0.0420 0.0991
RDFlex Lasso -0.0011 0.0069 0.9520 0.0288 0.0987
RDFlex Stacking -0.0014 0.0069 0.9480 0.0282 0.0989
2D Euclidean Fuzzy RDD Conventional Covs -0.0226 0.0119 0.4360
RDD Without Covs -0.0229 0.0122 0.4640
RDFlex Boosting -0.0234 0.0094 0.4000 0.0255 0.0744 0.1054 0.0654
RDFlex Global Lasso -0.0208 0.0106 0.5920 0.0349 0.1123 0.1037 0.2470
RDFlex Lasso -0.0226 0.0094 0.4320 0.0250 0.4624 0.1020 0.4017
RDFlex Stacking -0.0223 0.0098 0.4640 0.0252 0.0179 0.1025 0.0154
2D Euclidean Sharp RDD Conventional Covs -0.0255 0.0119 0.3120
RDD Without Covs -0.0257 0.0122 0.3160
RDFlex Boosting -0.0262 0.0104 0.3560 0.0254 0.1068
RDFlex Global Lasso -0.0253 0.0102 0.3440 0.0350 0.1038
RDFlex Lasso -0.0257 0.0103 0.3480 0.0247 0.1031
RDFlex Stacking -0.0258 0.0103 0.3200 0.0252 0.1034
Table 6: Mean Bias and standard error, coverage and first stage prediction quality of different estimators in estimation with two-dimensional estimators.
B.2 Simulation for Acknowledging Operator
Cutoff $c_{D}$
Mean Bias s.e. Coverage RMSE left Log loss left RMSE right Log loss right
setting method
Fuzzy RDD Conventional Covs -0.0019 0.0102 0.9735
RDD Without Covs -0.0017 0.0103 0.9823
RDFlex Boosting -0.0018 0.0086 0.9912 0.0243 0.0580 0.0988 0.3837
RDFlex Global Lasso -0.0020 0.0086 1.0000 0.0395 0.1020 0.0964 0.5188
RDFlex Lasso -0.0021 0.0083 0.9823 0.0242 0.1130 0.0958 0.4996
RDFlex Stacking -0.0018 0.0085 0.9823 0.0243 0.0135 0.0960 0.3813
Sharp RDD Conventional Covs -0.0042 0.0067 0.9320
RDD Without Covs -0.0042 0.0067 0.9360
RDFlex Boosting -0.0042 0.0058 0.9200 0.0243 0.0984
RDFlex Global Lasso -0.0042 0.0057 0.9360 0.0396 0.0956
RDFlex Lasso -0.0042 0.0057 0.9360 0.0241 0.0951
RDFlex Stacking -0.0042 0.0057 0.9320 0.0243 0.0954
Sharp Subset RDD Conventional Covs 0.0010 0.0101 0.9720
RDD Without Covs 0.0009 0.0102 0.9720
RDFlex Boosting 0.0005 0.0089 0.9760 0.0261 0.1122
RDFlex Global Lasso 0.0009 0.0084 0.9720 0.0255 0.1083
RDFlex Lasso 0.0009 0.0085 0.9720 0.0247 0.1078
RDFlex Stacking 0.0009 0.0085 0.9720 0.0258 0.1080
Table 7: Mean Bias and standard error, coverage and first stage prediction quality of different estimators in estimation at $c_{D}$.
Cutoff $c_{Y}$
Mean Bias s.e. Coverage RMSE left Log loss left RMSE right Log loss right
setting method
Fuzzy RDD Conventional Covs -0.0160 0.0125 0.6267
RDD Without Covs -0.0193 0.0115 0.5067
RDFlex Boosting -0.0390 0.4699 0.9800 0.0265 0.0338 0.0784 0.3830
RDFlex Global Lasso 0.0835 3.5647 0.9867 0.0380 0.0134 0.0747 0.5802
RDFlex Lasso -0.0697 4.5786 0.9867 0.0276 0.0213 0.0744 0.4486
RDFlex Stacking 0.0001 1.7755 0.9600 0.0257 0.0096 0.0777 0.3278
Sharp RDD Conventional Covs -0.0035 0.0081 0.8480
RDD Without Covs -0.0044 0.0079 0.8600
RDFlex Boosting -0.0037 0.0069 0.8480 0.0257 0.0789
RDFlex Global Lasso -0.0040 0.0081 0.8600 0.0357 0.0768
RDFlex Lasso -0.0041 0.0086 0.8680 0.0271 0.0766
RDFlex Stacking -0.0042 0.0077 0.8600 0.0256 0.0771
Sharp Subset RDD Conventional Covs -0.0008 0.0081 0.9760
RDD Without Covs -0.0008 0.0080 0.9760
RDFlex Boosting -0.0007 0.0059 0.9720 0.0366 0.0985
RDFlex Global Lasso -0.0006 0.0056 0.9800 0.0339 0.0970
RDFlex Lasso -0.0006 0.0056 0.9800 0.0336 0.0971
RDFlex Stacking -0.0007 0.0055 0.9760 0.0333 0.0970
Table 8: Mean Bias and standard error, coverage and first stage prediction quality of different estimators in estimation at $c_{Y}$.
Acknowledgements
This work was funded by the Bavarian Joint Research Program (BayVFP) – Digitization (Funding reference: DIK0294/01). The research partners ams OSRAM and Economic AI kindly thank the the VDI/VDE-IT Munich for the organization and the Free State of Bavaria for the financial support.